### v1

# Lie ideals in properly infinite C*-algebras

Abstract.
We show that every Lie ideal in a unital, properly infinite $C^{*}$-algebra is commutator equivalent to a unique two-sided ideal. It follows that the Lie ideal structure of such a $C^{*}$-algebra is concisely encoded by its lattice of two-sided ideals. This answers a question of Robert in this setting.
We obtain similar structure results for Lie ideals in unital, real rank zero $C^{*}$-algebras without characters. As an application, we show that every Lie ideal in a von Neumann algebra is related to a unique two-sided ideal, which solves a problem of Brešar, Kissin, and Shulman.
Key words and phrases:
Lie ideals, $C^{*}$-algebras, properly infinite, von Neumann algebras, commutators, square-zero elements
2020 Mathematics Subject Classification:
Primary 46L05, 46L10. Secondary 16W10, 17B60, 47B47.
The first named author was partially supported by the Knut and Alice Wallenberg Foundation (KAW 2021.0140).
1. Introduction
A Lie ideal in a $C^{*}$-algebra $A$ is a linear subspace $L\subseteq A$ such that for every $a\in A$ and $x\in L$ the commutator $[a,x]:=ax-xa$ belongs to $L$. The study of Lie ideals in $C^{*}$-algebras has a long history, and we refer to Section 2 in [Mar10] for an overview of available results and directions of active research.
Immediate examples of Lie ideals in a $C^{*}$-algebra $A$ are all subspaces contained in the center $Z(A)$ and all subspaces containing the commutator subspace $[A,A]$. If $A$ is unital and simple, then this does in fact describes all Lie ideals – a consequence of Herstein’s description of Lie ideals in simple rings [Her55]. Since the center of a unital, simple $C^{*}$-algebra $A$ contains only scalar multiples of the identity, a subspace $L$ is a Lie ideal in $A$ if and only if $L=\{0\}$, $L={\mathbb{C}}1$, or $[A,A]\subseteq L$.
Pop [Pop02] showed that a unital $C^{*}$-algebra satisfies $A=[A,A]$ if and only if $A$ has no tracial states. Consequently, a simple, unital $C^{*}$-algebra $A$ without tracial states has exactly three Lie ideals: $\{0\}$, ${\mathbb{C}}1$ and $A$. Cuntz and Pedersen [CP79] gave a description of the closure $\overline{[A,A]}$ as the intersection of the kernels of all tracial states on $A$. In particular, a simple, unital $C^{*}$-algebra $A$ with a unique tracial state has exactly four closed Lie ideals: $\{0\}$, ${\mathbb{C}}1$, $\overline{[A,A]}$ and $A$. This was first observed by Marcoux and Murphy [MM98], generalizing the case of UHF-algebras, which was first proved by Marcoux [Mar95].
Ng and Robert [NR16] showed that $[A,A]$ is closed in every pure, exact $C^{*}$-algebra. It follows that a simple, unital, exact, pure $C^{*}$-algebra $A$ with a unique tracial state has exactly four Lie ideals: $\{0\}$, ${\mathbb{C}}1$, $[A,A]$ and $A$. This applies for example to UHF-algebras, to irrational rotation algebras, and to the Jiang-Su algebra (where it was first shown by Ng [Ng12]). It also applies to many, and possibly all, simple, reduced group $C^{*}$-algebras. Indeed, if $G$ is a countable, exact, acylindrically hyperbolic group with trivial finite radical and with the rapid decay property, then $C^{*}_{\mathrm{red}}(G)$ has strict comparison [AGEP24], and therefore is pure [APTV24], and has a unique tracial state [BKKO17].
Turning towards Lie ideals in non-simple $C^{*}$-algebras, it is clear that the structure of the lattice of two-sided ideals will play a role. A systematic study of the interplay between Lie and two-sided ideals was initiated by Brešar, Kissin, and Shulman [BKS08] We first introduce the necessary notation. Following the usual convention, given subspaces $M,N\subseteq A$, we use $[M,N]$ to denote the subspace of $A$ generated by the set of commutators $[x,y]$ for $x\in M$ and $y\in N$. Further, given a subspace $K\subseteq A$, we set
$T(K):=\big{\{}a\in A:[A,a]\subseteq K\big{\}}.$
This ‘derived subspace’ was first considered by Herstein [Her55], and it is denoted by $N(K)$ in [BKS08].
Given a two-sided ideal $I\subseteq A$, Brešar, Kissin, and Shulman observed that a subspace $L\subseteq A$ is a Lie ideal whenever
$[A,I]\subseteq L\subseteq T([A,I]).$
We say that a Lie ideal $L$ is embraced by $I$ if it satisfies these inclusions. The Lie ideals that are embraced by some two-sided ideal are thought of as ‘unsurprising’. A major question is if there exist any Lie ideals that do not arise this way:
[…]
Given a $C^{*}$-algebra $A$, is every Lie ideal of $A$ embraced by some two-sided ideal of $A$?
To study this problem, two further notions were introduced in [BKS08]: Given a Lie ideal $L\subseteq A$ and a two-sided ideal $I\subseteq A$, we say that $L$ is commutator equivalent to $I$ if
$[A,I]=[A,L].$
Further, we say that $L$ is related to $I$ if
$[A,I]\subseteq L\subseteq T(I).$
It is easy to see that these constitute a sufficient and a necessary conditions for embracement. Indeed, if $L$ is commutator equivalent to $I$, then $L$ is embraced by $I$. Further, if $L$ is embraced by $I$, then $L$ is related to $I$.
It was shown in [BKS08, Theorem 5.27] that every Lie ideal $L$ in a $C^{*}$-algebra $A$ is topologically commutator equivalent to a two-sided ideal $I$ in the sense that $[A,L]$ and $[A,I]$ have the same closure. This leads to a description of all closed Lie ideals in terms of closed two-sided ideals. However, this does not show that all closed Lie ideals in a $C^{*}$-algebra are commutator equivalent to (let alone embraced by or related to) a two-sided ideal. For general non-closed Lie ideals even less is known, and Section 1 as well as the following stronger problem remain open:
[…]
Given a $C^{*}$-algebra $A$, is every Lie ideal of $A$ commutator equivalent to some two-sided ideal of $A$?
The results mentioned at the beginning of the introduction show that every Lie ideal in a unital, simple $C^{*}$-algebra $A$ is commutator equivalent to a two-sided ideal, which gives positive solutions to Problems 1 and 1 in this setting. Indeed, the only two-sided ideals are $\{0\}$ and $A$, and a Lie ideal is commutator equivalent to $\{0\}$ if and only if it is contained in $Z(A)$, while a Lie ideal it is commutator equivalent to $A$ if and only if it contains $[A,A]$. (This uses that $[A,A]=[A,[A,A]]$, which follows for example from [GT24, Corollary 3.4].)
In [FMS82], Fong, Miers and Sourour showed that every Lie ideal in the von Neumann algebra of bounded, linear operators on a separable Hilbert space is embraced by a two-sided ideal. This was substantially extended by Brešar, Kissin, and Shulman, who showed in [BKS08, Theorem 5.19] that every Lie ideal in a von Neumann algebra is commutator equivalent to a two-sided ideal, thereby providing positive solutions to Problems 1 and 1 for von Neumann algebras.
Our main result is a positive solution to Section 1 for unital, properly infinite $C^{*}$-algebras.
[…]
Let $L$ be a Lie ideal in a unital, properly infinite $C^{*}$-algebra $A$. Then $L$ is commutator equivalent to the two-sided ideal $I:=A[A,L]A$, and hence also embraced by $I$, and related to $I$. Further, $I$ is the only two-sided ideal of $A$ to which $L$ is related.
The proof of Section 1 is inspired by a result of Marcoux [Mar95], who showed that if $B$ is a unital $C^{*}$-algebra and $n\geq 2$, then every Lie ideal in the matrix algebra $A=M_{n}(B)$ is related to a two-sided ideal in $A$. Our main technical advancement is a generalization of this result to unital $C^{*}$-algebras $A$ that admit a unital *-homomorphism $M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})\to A$.
The next result summarizes Theorems 2, 2, 3, and 3
[…]
Let $A$ be a unital $C^{*}$-algebra that admits a unital *-homomorphism $M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})\to A$. The following statements hold:
  1. (1)
    Given a Lie ideal $L\subseteq A$, for the two-sided ideal $I=A[A,L]A$, we have
    $[A,I]=[A,[A,L]]\subseteq[A,L]\subseteq I,\,\,\,\text{ and }\,\,\,[A,I]\subseteq L\subseteq T(I).$
    Further, $I$ is the only two-sided ideal of $A$ to which $L$ is related.
  2. (2)
    A Lie ideal $L\subseteq A$ is embraced by some two-sided ideal (which then necessarily is $A[A,L]A$) if and only if $L$ is commutator equivalent to some two-sided ideal, if and only if $[A,L]=[A,[A,L]]$.
  3. (3)
    Given a two-sided ideal $I\subseteq A$, we have
    $I=A[A,I]A,\,\,\,\text{ and }\,\,\,[A,I]=[A,[A,I]].$
  4. (4)
    There is a natural bijection between the set $\mathcal{I}$ of two-sided ideals in $A$ and the set
    =$\mathcal{L}=\big{\{}L\subseteq A\text{ Lie ideal}:L=[A,L]\big{\}}.$
    given by the maps $I\mapsto[A,I]$ and $L\mapsto A[A,L]A$.
We note that Section 1 applies to all unital, properly infinite $C^{*}$-algebras, which then yields Section 1. It also applies to unital, real rank zero $C^{*}$-algebras without characters, and in particular to all von Neumann algebras with zero commutative summand, which we use to recover the solution to Section 1 for von Neumann algebras from [BKS08]. Our approach also shows that a Lie ideal in a von Neumann algebra without commutative summand is related to a unique two-sided ideal, which solves [BKS08, Problem 5.21].
[…]
Let $L$ be a Lie ideal in a von Neumann algebra $M$. Then $L$ is commutator equivalent to the two-sided ideal $I:=M[M,L]M$, and hence also embraced by $I$, and related to $I$.
If $M$ has zero commutative summand, then $I$ is the only two-sided ideal of $M$ to which $L$ is related.
Conventions
Given a $C^{*}$-algebra, we use $A_{+}$ to denote its collection of positive elements, and we use $\widetilde{A}$ to denote the minimal unitization. The absolute value of an element $x\in A$ is defined as $|x|:=(x^{*}x)^{\frac{1}{2}}$.
Acknowledgements
The author thanks Leonel Robert for valuable comments on the Dixmier property for $C^{*}$-algebras, and Matej Brešar for useful feedback on an earlier version of this paper.
2. Lie ideals in C*-algebras with weakly divisible unit
In this section, we prove that every Lie ideal $L$ in a $C^{*}$-algebra $A$ that admits a unital *-homomorphism $M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})\to A$ is related to a unique two-sided ideal $I$ in $A$; see Section 2. We further show that $L$ is commutator equivalent to $I$ if and only if $[A,L]=[A,[A,L]]$; see Section 2.
Following [PR04, Definition 5.1], a non-zero projection $p$ in a $C^{*}$-algebra is said to be weakly divisible of degree $n$ if there exists a unital *-homomorphism
$M_{n_{1}}({\mathbb{C}})\oplus M_{n_{2}}({\mathbb{C}})\oplus\ldots\oplus M_{n_{r}}({\mathbb{C}})\to pAp$
for some natural numbers $r\geq 1$ and $n_{1},\ldots,n_{r}\geq n$. The *-homomorphism may additionally be assumed to be injective.
As noted in [PR04, p.164], every matrix algebra $M_{n}({\mathbb{C}})$ for $n\geq 2$ admits a unital (not necessarily injective) *-homomorphism $M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})\to M_{n}({\mathbb{C}})$, and it follows that a projection $p$ weakly divisible of degree $2$ if and only if there exists a unital (not necessarily injective) *-homomorphism $M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})\to pAp$. We deduce that the unit in a $C^{*}$-algebra $A$ is weakly divisible of degree $2$ if and only if there exists an injective *-homomorphism
$M_{2}({\mathbb{C}})\to A,\quad M_{3}({\mathbb{C}})\to A,\quad\text{ or }\quad M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})\to A.$
In the first two cases, we get $A\cong M_{2}(B)$ and $A\cong M_{3}(C)$ for suitable unital $C^{*}$-algebras $B$ and $C$, and for such matrix algebras it was shown by Marcoux [Mar95, Theorem 2.6] that every Lie ideal is related to a two-sided ideal (which is also unique, as shown in [BKS08, Corollary 4.18]).
Our main contribution to Section 2 is therefore in the case that $A$ admits a unital (injective) *-homomorphism $M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})\to A$ but no unital *-homomorphism $M_{n}({\mathbb{C}})\to A$ for $n=2$ or $n=3$ (or actually any $n\geq 2$). The Cuntz algebra $\mathcal{O}_{\infty}$ is an important such example.
Our approach is inspired by the methods developed by Marcoux [Mar95] to describe Lie ideals in matrix algebras, as well as the techniques of Brešar, Kissin, and Shulman [BKS08, Section 3] to study Lie ideals in algebras that decompose with respect to an idempotent. However, the result does not simply follow by combining these methods, for example by applying [BKS08, Theorem 3.4] for a suitable idempotent in the $C^{*}$-algebra. Specifically, given a unital *-homomorphism $M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})\to A$, we do not decompose $A$ with respect to the projection $p$ given by the image of the unit of $M_{2}({\mathbb{C}})$, in which case $pAp+(1-p)A(1-p)$ is a direct sum of matrix algebras (whence its Lie ideals are related to two-sided ideals), but it is unclear if (3.17) in [BKS08, Theorem 3.4] holds (or if $p$ is locally cyclic in the sense of [BKS08, Definition 3.5]). Instead, we use a ‘mixed’ decomposition with respect to a projection $g$ that is given by the image of the sum of a rank-one projection in $M_{2}({\mathbb{C}})$ with a rank-one projection in $M_{3}({\mathbb{C}})$, as explained in Section 2. In this case, one can show that $g$ is locally cyclic, which is the idea underlying the proof that $A_{12}LA_{12}\subseteq[A,[A,L]]$ in Lemma 2.1. On the other hand, the algebra $gAg+(1-g)A(1-g)$ is not a sum of matrix algebras and it is therefore not immediate that its Lie ideals are related to two-sided ideals.
[…]
Let $A$ be a unital $C^{*}$-algebra that admits a unital *-homomorphism
$\varphi\colon M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})\to A.$
Let $(e_{ij})_{i,j=1,2}$ denote the image of matrix units in $M_{2}({\mathbb{C}})$ under $\varphi$, and similarly let $(f_{ij})_{i,j=1,2,3}$ denote the image of matrix units in $M_{3}({\mathbb{C}})$ under $\varphi$. We do not assume that $\varphi$ is injective, and therefore the $e_{ij}$ or $f_{ij}$ may be zero.
Set
$g:=e_{11}+f_{11},\,\,\,\text{ and }\,\,\,h:=e_{22}+f_{22}+f_{33}.$
Then $1=g+h$, and we view elements in $A$ as operator matrices with respect to the decomposition induces by $g$ and $h$. More formally, we set
$A_{11}:=gAg,\quad A_{12}:=gAh,\quad A_{21}:=hAg,\,\,\,\text{ and }\,\,\,A_{22}:=hAh,$
and every element $a$ in $A$ can be written uniquely as a sum $a=a_{11}+a_{12}+a_{21}+a_{22}$ with $a_{ij}\in A_{ij}$ for $i,j=1,2$.
Lemma 2.1.
Retain the notation from Section 2, and let $L\subseteq A$ be a Lie ideal. Then the sets
$A_{11}LA_{22},\quad A_{22}LA_{11},\quad A_{12}LA_{12},\,\,\,\text{ and }\,\,\,A_{21}LA_{21}$
are all contained in $[A,[A,L]]$, and thus also in $[A,L]$ and in $L$.
Proof.
We verify that $A_{11}LA_{22}\subseteq[A,[A,L]]$. We first show that $gLh\subseteq L$. To see this, let $x\in L$. Then
$gx-xg=[g,x]\in[A,L],\,\,\,\text{ and }\,\,\,gx-2gxg+xg=[g,[g,x]]\in[A,[A,L]]\subseteq[A,L].$
Adding these expressions, we get
$2gxh=2gx-2gxg\in[A,L],$
and using that $hg=0$ we get
$gxh=[\tfrac{1}{2}g,gxh]\in[A,L],$
We have shown that $gLh\subseteq L$.
Now, given $x\in L$ and $a,b\in A$, we have
$gagxhbh=[[gag,gxh],hbh]\in[[A,L],A]=[A,[A,L]].$
We have shown that $A_{11}LA_{22}\subseteq[A,[A,L]]$.
Analogously, one verifies that $A_{22}LA_{11}\subseteq[A,[A,L]]$.
We verify that $A_{12}LA_{12}\subseteq[A,[A,L]]$. Given $v\in A_{12}$ and $y\in L$, we have
(1) $vxv=[-\tfrac{1}{2}v,[v,y]]\in[A,[A,L]].$
Set
$v_{1}:=e_{12}+f_{12},\quad w_{1}:=e_{21}+f_{21},\quad v_{2}:=e_{12}+f_{13},\,\,\,\text{ and }\,\,\,w_{2}:=e_{21}+f_{31}.$
We have
$v_{1}w_{1}=v_{2}w_{2}=e_{11}+f_{11}=g,\quad w_{1}v_{1}=e_{22}+f_{22},\,\,\,\text{ and }\,\,\,w_{2}v_{2}=e_{22}+f_{33}.$
Now, given $x\in L$ and $a,b\in A$, set
$b^{\prime}=b(\tfrac{1}{2}e_{22}+f_{22}+f_{33}).$
Then $b^{\prime}$ belongs to $gAh$ and we have
$b=b^{\prime}(2e_{22}+f_{22}+f_{33}).$
We have already proved that $gAgLhAh\subseteq L$, which gives
$w_{1}axb^{\prime}w_{1}\in L,\,\,\,\text{ and }\,\,\,w_{2}axb^{\prime}w_{2}\in L.$
Applying (1) for $y=w_{1}axb^{\prime}w_{1}$ and $v=v_{1}$, we get
$axb^{\prime}(e_{22}+f_{22})=v_{1}(w_{1}axb^{\prime}w_{1})v_{1}\in[A,[A,L]].$
Analogously, we obtain that
$axb^{\prime}(e_{22}+f_{33})=v_{2}(w_{2}axb^{\prime}w_{2})v_{2}\in[A,[A,L]],$
and consequently
$axb=axb^{\prime}(2e_{22}+f_{22}+f_{33})\in[A,[A,L]].$
We have shown that $A_{12}LA_{12}\subseteq[A,[A,L]]$.
Analogously, one verifies that $A_{21}LA_{21}\subseteq[A,[A,L]]$. ∎
Lemma 2.2.
Retain the notation from Section 2, and let $L\subseteq A$ be a Lie ideal. Then
$AgLhA=AhLgA.$
Proof.
Using at the first step that $A=AgA=AhA$ since $g$ and $h$ are full, and applying at the third step that $hAgLhAg\subseteq L$ by Lemma 2.1, we get
$AgLhA=AhAgLhAgA=Ah(hAgLhAg)gA\subseteq AhLgA.$
Analogously, we have
$AhLgA=AgAhLgAhA=Ag(gAhLgAh)hA\subseteq AgLhA,$
as desired. ∎
Lemma 2.3.
Retain the notation from Section 2, and let $L\subseteq A$ be a Lie ideal. Then
$[A,L]\subseteq AgLhA.$
Proof.
To simplify notation, we set $I=AgLhA$. We have
$[A,gLh]\subseteq AgLh+gLhA\subseteq I.$
Further, using Lemma 2.2 at the third step, we have
$[A,hLg]\subseteq AhLg+hLgA\subseteq AhLgA=AgLhA=I.$
Thus, given $x\in L$, we have $[A,gxh+hxg]\subseteq I$. Note that $gxh,hxg\in L$ by Lemma 2.1. It follows that
$gxg+hxh=x-gxh+hxg\in L.$
We need to show that $[A,gxg+hxh]\subseteq I$. Using additivity of the Lie bracket and that $I$ is an additive subgroup, it suffices to verify that $[A_{ij},gxg+hxh]\subseteq I$ for $i,j=1,2$.
We show that $[A_{12},gxg+hxh]\subseteq I$. Given $a\in A$, we have
$[gah,gxg+hxh]=gahxh-gxgah.$
Using that $gxg+hxh\in L$, we have $[gah,gxg+hxh]\in L$, and thus
$[gah,gxg+hxh]\in L\cap gAh=gLh\subseteq AgLhA=I.$
We have shown that $[A_{12},gxg+hxh]\subseteq I$.
Similarly, and using also Lemma 2.2 at the last step, we get
$[hag,gxg+hxh]=hagxg-hxhag\in L\cap hAg=hLg\subseteq AhLgA=I$
for every $a\in A$. This shows that $[A_{21},gxg+hxh]\subseteq I$.
We show that $[A_{11},gxg+hxh]\subseteq I$. Given $a\in A$, set $y=gagxg-gxgag$. Then
$y=gagxg-gxgag=[gag,gxg+hxh]\in[A,L]\subseteq L.$
Then
$y(e_{12}+f_{12})=[y,e_{12}+f_{12}]\in[L,A]\subseteq L$
and thus
$y(e_{12}+f_{12})\in L\cap gAh=gLh\subseteq I.$
We deduce that
$y=y(e_{12}+f_{12})(e_{21}+f_{21})\in IA=I.$
We show that $[A_{22},gxg+hxh]\subseteq I$. Given $b\in A$, set $z=hbhxh-hxhbh$. Then
$z=hbhxh-hxhbh=[hbh,gxg+hxh]\in L\cap A_{22}.$
Then
$(e_{12}+f_{12})z=[e_{12}+f_{12},z]\in L\cap gAh\subseteq I,\,\,\,\text{ and }\,\,\,f_{13}z=[f_{13},z]\in I.$
We deduce that
$z=(e_{21}+f_{21})(e_{12}+f_{12})z+f_{31}f_{13}z\in AI+AI=I,$
as desired. ∎
[…]
Retain the notation from Section 2, and let $L\subseteq A$ be a Lie ideal. Then
$A[A,L]=A[A,L]A=AgLhA=AhLgA.$
Proof.
In general, if $M\subseteq A$ is a Lie ideal, then $AMA=AM$; see, for example, [GLT24, Lemma 3.2]. Since $[A,L]$ is a Lie ideal, we get $A[A,L]=A[A,L]A$.
By Lemma 2.3, we have $[A,L]\subseteq AgLhA$, and thus $A[A,L]A\subseteq AgLhA$. By Lemma 2.2, we have $AgLhA=AhLgA$. Finally, by Lemma 2.1, we have $gLh\subseteq[A,L]$, and thus $AgLhA\subseteq A[A,L]A$. ∎
Lemma 2.4.
Retain the notation from Section 2, and let $L\subseteq A$ be a Lie ideal, and set $I=A[A,L]A$. Then
$[A,I]\subseteq[A,[A,L]].$
Proof.
For each $i,j=1,2$, set $I_{ij}:=I\cap A_{ij}$. Since $I$ is a two-sided ideal, we have $I=I_{11}+I_{12}+I_{21}+I_{22}$. By Section 2, we have $I=AgLhA$, and thus
$I=AgLhA=\underbrace{gAgLhAg}_{I_{11}}+\underbrace{gAgLhAh}_{I_{12}}+\underbrace{hAgLhAg}_{I_{21}}+\underbrace{hAgLhAh}_{I_{22}}.$
By additivity of the Lie bracket and since $[A,L]$ is an additive subgroup, it suffices to show that $[A,I_{ij}]\subseteq[A,[A,L]]$ for each $i,j=1,2$.
By Lemma 2.1, we have
$I_{12}=gAgLhAh\subseteq[A,L],\,\,\,\text{ and }\,\,\,I_{21}=hAgLhAg\subseteq[A,L],$
and thus $[A,I_{12}]\subseteq[A,[A,L]]$ and $[A,I_{21}]\subseteq[A,[A,L]]$.
We show that $[A,I_{11}]\subseteq[A,[A,L]]$. It suffices to verify that $[A_{ij},I_{11}]\subseteq[A,[A,L]]$ for each $i,j=1,2$. Using that $hg=0$, and using Lemma 2.1 at the last step, we get
$[A_{12},I_{11}]=[gAh,gAgLhAg]=-(gAgLhAg)(gAh)=gAgLhAgAh\subseteq[A,[A,L]].$
Similarly, we have
$[A_{21},I_{11}]=[hAg,gAgLhAg]=(hAg)(gAgLhAg)=hAgAgLhAg\subseteq[A,[A,L]].$
We also have $[A_{22},I_{11}]=\{0\}\subseteq[A,[A,L]]$.
Further, using Lemma 2.1 at the last step, we have
$I_{11}=(gAgLh)Ag\subseteq[gAgLh,Ag]+Ag(gAgLh)\subseteq[A,L]+AgLh.$
Using again Lemma 2.1 at the last step, we get
$[A_{11},I_{11}]$ $\subseteq[A_{11},[A,L]]+[A_{11},AgLh]\subseteq[A,[A,L]]+[gAg,AgLh]$
$\subseteq[A,[A,L]]+gAgAgLh\subseteq[A,[A,L]].$
We have verified that $[A,I_{11}]\subseteq[A,[A,L]]$.
Analogously, one shows that $[A,I_{22}]\subseteq[A,[A,L]]$, which completes the proof. ∎
The next result verifies statement (1) in Section 1.
[…]
Let $A$ be a unital $C^{*}$-algebra that admits a unital *-homomorphism $M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})\to A$, and let $L\subseteq A$ be a (not necessarily closed) Lie ideal. Then, for the two-sided ideal $I:=A[A,L]A$, we have
(2) $[A,I]=[A,[A,L]]\subseteq[A,L]\subseteq I,\,\,\,\text{ and }\,\,\,[A,I]\subseteq L\subseteq T(I).$
Further, $I$ is the only two-sided ideal of $A$ to which $L$ is related.
Proof.
Using Lemma 2.3 at the first step, and Section 2 at the second step, we have
$[A,L]\subseteq AgLhA=I.$
This implies that $[A,[A,L]]\subseteq[A,I]$, and the converse inclusion $[A,I]\subseteq[A,[A,L]]$ is shown in Lemma 2.4. Using that $L$ is a Lie ideal, we have $[A,L]\subseteq L$, and therefore $[A,[A,L]]\subseteq[A,L]$. This shows the left chain of inclusions in (2).
The inclusion $[A,I]\subseteq L$ follows using that $[A,I]\subseteq[A,L]$ and $[A,L]\subseteq L$. Finally, since $[A,L]\subseteq I$, we have $L\subseteq T(I)$.
To show uniqueness of $I$, let $J$ be a two-sided ideal that is related to $L$, that is, such that
$[A,J]\subseteq L\subseteq T(J).$
Then $[A,L]\subseteq J$, and since $I$ is the two-sided ideal generated by $[A,L]$ we get $I\subseteq J$.
Applying Lemma 2.1 at the first step for $J$, considered as a Lie ideal, we get
$gJh\subseteq[A,J]\subseteq L,$
and thus
$gJh\subseteq L\cap gAh=gLh.$
Since $g$ and $h$ are full, we have $A=AgA$ and $A=AhA$, and consequently
$J=AJA=AgAJAhA=AgJhA\subseteq AgLhA=I.$
In conclusion, we get $I=J$, as desired. ∎
The next result verifies statement (2) in Section 1.
[…]
Let $A$ be a unital $C^{*}$-algebra that admits a unital *-homomorphism $M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})\to A$, and let $L\subseteq A$ be a (not necessarily closed) Lie ideal. Then the following statements are equivalent:
  1. (1)
    The Lie ideal $L$ is commutator equivalent to the two-sided ideal $A[A,L]A$.
  2. (2)
    The Lie ideal $L$ is embraced by some two-sided ideal.
  3. (3)
    We have $[A,L]=[A,[A,L]]$.
Proof.
It is clear that (1) implies (2). Assuming (2), let us show that (3) holds. Set $I:=A[A,L]A$, and let $J\subseteq A$ be some two-sided ideal that embraces $L$. Then $L$ is also related to $J$. It follows that $I=J$, since $I$ is the unique two-sided ideal to which $L$ is related, by Section 2. It follows that $L\subseteq T([A,I])$. Using this at the first step, and using Section 2 at the second step, we get
$[A,L]\subseteq[A,I]=[A,[A,L]],$
which verifies (3).
Assuming (3), let us show that (1) holds. Using that assumption at the firs step, and Section 2 at the second step, we get
$[A,L]=[A,[A,L]]=[A,I],$
which verifies (1). ∎
The next result is essentially contained in [PR04, Proposition 5.7]. We only show how to remove the separability assumption. A character on a $C^{*}$-algebra $A$ is one-dimensional, irreducible representation, that is, a surjective *-homomorphism $A\to{\mathbb{C}}$.
[…]
Let $A$ be a unital $C^{*}$-algebra of real rank zero. Then $A$ has no characters if and only if $A$ admits a unital *-homomorphism $M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})\to A$.
Proof.
The backwards implication is clear. To show the forward implication, assume that $A$ has no characters. By [KR15, Lemma 3.5], there exists a separable sub-$C^{*}$-algebra $A_{0}\subseteq A$ containing the unit of $A$. Using that real rank zero is separably inheritable (see [Bla06, Section II.8.5]), we obtain a separable sub-$C^{*}$-algebra $B\subseteq A$ of real rank zero with $A_{0}\subseteq B$. It follows that $B$ has no characters, which allows us to apply [PR04, Proposition 5.7] for $B$. We obtain a unital *-homomorphism from $M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})$ to $B$, and thus also to $A$. ∎
[…]
Let $A$ be a unital $C^{*}$-algebra of real rank zero that has no characters. Then Section 2 applies to $A$. In particular, every Lie ideal in $A$ is related to a (unique) two-sided ideal.
We expect that the following question has a positive answer.
[…]
Let $A$ be a unital $C^{*}$-algebra that has no characters. Is every Lie ideal in $A$ is related to a (unique) two-sided ideal?
3. Two-sided ideals in C*-algebras with weakly divisible unit
In this section, we prove that every two-sided ideal $I$ in a $C^{*}$-algebra $A$ that admits a unital *-homomorphism $M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})\to A$ satisfies $I=A[A,I]A$, and we describe $[A,I]$ as the subspace spanned by certain square-zero elements in $I$; see Section 3. If $I$ is a Dixmier ideal (in which case there is a well-defined ideal $I^{\frac{1}{2}}$ satisfying $(I^{\frac{1}{2}})^{2}=I$), then $I=[A,I]+[A,I^{\frac{1}{2}}]^{2}$; see Section 3. In this case every element in $I$ is a sum of square-zero elements in $I$ and of products of pairs of square-zero elements in $I^{\frac{1}{2}}$.
The next definition generalizes the notion of orthogonally factorizable square-zero elements in a ring ([GT23, Definition 5.1]) to a relative notion with respect to a two-sided ideal.
[…]
Let $I$ be a two-sided ideal in a $C^{*}$-algebra $A$. We let $N_{2}(I)$ denote the square-zero elements in $I$, that is,
$\operatorname{N}_{2}(I):=\big{\{}x\in I:x^{2}=0\big{\}}.$
We further set
$\operatorname{FN}_{2}(A,I):=\big{\{}x\in I:\text{ there exist }a,b\in A,y\in I\text{ with }x=ayb\mbox{ and }ba=0\big{\}}.$
Lemma 3.1.
Let $I\subseteq A$ be a two-sided ideal in a $C^{*}$-algebra, and let $x\in\operatorname{FN}_{2}(A,I)$. Then there exist $a,b\in A_{+}$ and $y\in\operatorname{FN}_{2}(A,I)\subseteq I$ such that
$x=ayb,\,\,\,\text{ and }\,\,\,ab=ba=0.$
Proof.
By definition, there are $c,d\in A$ and $z\in I$ such that
$x=czd,\,\,\,\text{ and }\,\,\,dc=0.$
Let $c=v|c|$ and $d=w|d|$ the polar decompositions in $A^{**}$ and set
$a:=|c^{*}|^{\frac{1}{2}},\quad b:=|d|^{\frac{1}{2}},\,\,\,\text{ and }\,\,\,y:=|c^{*}|^{\frac{1}{2}}vxw|d|^{\frac{1}{2}}.$
By properties of the polar decomposition (see, for example, [GKT23, Proposition 2.1]), we have $|c^{*}|^{\frac{1}{2}}v,w|d|^{\frac{1}{2}}\in A$, and thus $y\in A$. Further, we have $x=ayb$.
Since $dc=0$, we have $(d^{*}d)^{n}(cc^{*})^{m}=0$ for every $m,n\geq 1$. It follows that $p(d^{*}d)q(cc^{*})=0$ for every polynomials $p$ and $q$ with vanishing constant terms. Using functional calculus, we get $f(d^{*}d)g(cc^{*})=0$ for every continuous functions $f,g\colon\mathbb{R}\to\mathbb{R}$ with $f(0)=g(0)=0$. In particular, we get $|d|^{\frac{1}{2}}|c^{*}|^{\frac{1}{2}}=0$. Thus, we have $ba=0$, and then also $ab=0$.
To show that $y$ belongs to $\operatorname{FN}_{2}(A,I)$, consider the decomposition
$y=|c^{*}|^{\frac{1}{4}}\big{(}|c^{*}|^{\frac{1}{4}}vxw|d|^{\frac{1}{4}}\big{)}|d|^{\frac{1}{4}}.$
Then $|c^{*}|^{\frac{1}{4}}vxw|d|^{\frac{1}{4}}$ belongs to $I$, and we have also seen above that $|c^{*}|^{\frac{1}{4}}$ and $|d|^{\frac{1}{4}}$ have zero product, whence $y\in\operatorname{FN}_{2}(A,I)$. ∎
The next result is analogous to [GT23, Lemma 5.2].
Lemma 3.2.
Let $I\subseteq A$ be a two-sided ideal in a $C^{*}$-algebra. Then
$\operatorname{FN}_{2}(A,I)\subseteq[A,\operatorname{FN}_{2}(A,I)]\subseteq[A,[A,I]]\subseteq[A,I],\,\,\,\text{ and }\,\,\,\operatorname{FN}_{2}(A,I)\subseteq\operatorname{N}_{2}(I).$
Proof.
Let $x\in\operatorname{FN}_{2}(A,I)$. By Lemma 3.1, we can pick $a,b\in A_{+}$ and $y\in I$ such that $x=ayb$ and $ba=0$. As in the proof of Lemma 3.1, we see that $ba^{\frac{1}{2}}=0$. It follows that $a^{\frac{1}{2}}yb\in\operatorname{FN}_{2}(A,I)$ and then
$x=ayb=\big{[}a^{\frac{1}{2}},a^{\frac{1}{2}}yb\big{]}\in[A,\operatorname{FN}_{2}(A,I)].$
Since $\operatorname{FN}_{2}(A,I)$ is a subset of $I$, we have $[A,\operatorname{FN}_{2}(A,I)]\subseteq[A,I]$. Combining these results, and using at the last step that $[A,I]$ is a Lie ideal, we get
$\operatorname{FN}_{2}(A,I)\subseteq[A,\operatorname{FN}_{2}(A,I)]\subseteq[A,[A,\operatorname{FN}_{2}(A,I)]]\subseteq[A,[A,I]]\subseteq[A,I].$
Given $x=ayb\in\operatorname{FN}_{2}(A,I)$, with $a,b\in A$ and $y\in I$ such that $ba=0$, we have
$x^{2}=(ayb)(ayb)=0,$
and thus $x\in\operatorname{N}_{2}(I)$. ∎
[…]
We note that Section 3 and [GT23, Definition 5.1] agree when considering the ideal $I=A$ in a $C^{*}$-algebra $A$. Indeed, an element $x$ in $A$ belongs to $\operatorname{FN}_{2}(A)$ in the sense of [GT23, Definition 5.1] if there exist $s,t\in A$ such that $x=st$ and $ts=0$, while $x$ belongs to $\operatorname{FN}_{2}(A,A)$ as defined in Section 3 if there exist $a,y,b\in A$ such that $x=ayb$ and $ba=0$.
Given a factorization $x=ayb$ with $ba=0$, we can use $s=a$ and $t=yb$. Conversely, given a factorization $x=st$ with $ts=0$, it follows that $|t|^{\frac{1}{2}}s=0$. We consider the polar decomposition $t=v|t|$ in $A^{**}$. Then $v|t|^{\frac{1}{2}}$ belongs to $A$, and we can use $a=s$, $y=v|t|^{\frac{1}{2}}$ and $b=|t|^{\frac{1}{2}}$.
Given a subset $V$ in a complex vector space, we use $\operatorname{span}_{\mathbb{C}}V$ to denote the linear subspace generated by $V$. Following [GT23, Definition 2.6], we say that a ${\mathbb{C}}$-algebra $A$ is zero-product balanced if for all $a,b,c\in A$ we have
$ab\otimes c-a\otimes bc\in\operatorname{span}_{\mathbb{C}}\big{\{}u\otimes v\in A\otimes_{\mathbb{C}}A:uv=0\big{\}}\subseteq A\otimes_{\mathbb{C}}A.$
We see that the class of $C^{*}$-algebras considered in this section are zero-product balanced:
Lemma 3.3.
Let $A$ be a unital $C^{*}$-algebra that admits a unital *-homomorphism $M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})\to A$. Then $A$ is zero-product balanced as a ${\mathbb{C}}$-algebra.
Proof.
As in Section 2, let $(e_{ij})_{i,j=1,2}$ and $(f_{ij})_{i,j=1,2,3}$ be the images in $A$ of matrix units in $M_{2}({\mathbb{C}})$ and $M_{3}({\mathbb{C}})$. Then $e_{11}+f_{11}$ and $e_{22}+f_{22}$ are two full orthogonal projections in $A$. This allows us to apply [Rob16, Corollary 3.8], which shows that $A$ is generated by its projections as a ${\mathbb{C}}$-algebra, and thus is zero-product balanced by [GT23, Proposition 3.6]. ∎
A two-sided ideal $I\subseteq A$ in a $C^{*}$-algebra is said to be idempotent if $I=I^{2}$, that is, if every element in $I$ is of the form $x_{1}y_{1}+\ldots+x_{n}y_{n}$ for some $n\in{\mathbb{N}}$ and suitable $x_{j},y_{j}\in I$.
[…]
Let $A$ be a $C^{*}$-algebra that is zero-product balanced as a ${\mathbb{C}}$-algebra, and let $I\subseteq A$ be a two-sided ideal such that $I=AIA$ (for example, $A$ is unital or $I$ is idempotent). Then
$[A,I]=[A,[A,I]]=\operatorname{span}_{\mathbb{C}}\operatorname{FN}_{2}(A,I).$
Proof.
The inclusions $\supseteq$’ hold by Lemma 3.2. The proof of the converse inclusion is based on a combination of the methods used in the proofs of Proposition 2.14 and Theorem 5.3 in [GT23], which in turn are inspired by that of [Bre21, Theorem 9.1]. To verify that $[A,I]\subseteq\operatorname{span}_{\mathbb{C}}\operatorname{FN}_{2}(A,I)$, let $x\in A$ and $y\in I$. Set $F:=\operatorname{span}_{\mathbb{C}}\operatorname{FN}_{2}(A,I)$, and define
$\varphi_{y}\colon A\times A\to I/F,\quad\varphi_{y}(a,b):=bya+F.\qquad\qquad(a,b\in A)$
Note that $\varphi_{y}$ is a well-defined ${\mathbb{C}}$-bilinear map. Further, $\varphi_{y}$ preserves zero-products since if $a,b\in A$ satisfy $ab=0$, then $bya\in\operatorname{FN}_{2}(A,I)\subseteq F$. Using [GT23, Proposition 2.8] at the second step, we obtain that
$cyab+F=\varphi_{y}(ab,c)=\varphi_{y}(a,bc)=bcya+F$
and thus
$bcya-cyab\in F=\operatorname{span}_{\mathbb{C}}\operatorname{FN}_{2}(A,I)$
for all $a,b,c\in A$. The assumption that $I=AIA$ allows us to pick $n$ and $c_{j},a_{j}\in A$ and $y_{j}\in I$ for $j=1,\ldots,n$ such that $y=\sum_{j}c_{j}y_{j}a_{j}$. Then
$[x,y]=\sum_{j=1}^{n}[x,c_{j}y_{j}a_{j}]=\sum_{j=1}^{n}\big{(}xc_{j}y_{j}a_{j}-c_{j}y_{j}a_{j}x\big{)}\in\operatorname{span}_{\mathbb{C}}\operatorname{FN}_{2}(A,I),$
which implies that $[A,I]\subseteq\operatorname{span}_{\mathbb{C}}\operatorname{FN}_{2}(A,I)$. ∎
[…]
Does there exist a two-sided ideal $I$ in a $C^{*}$-algebra $A$ such that $[A,[A,I]]\neq[A,I]$?
The next result verifies statement (3) in Section 1.
[…]
Let $A$ be a unital $C^{*}$-algebra that admits a unital *-homomorphism $M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})\to A$, and let $I\subseteq A$ be a (not necessarily closed) two-sided ideal. Then
$I=A[A,I]A=A[A,I],\,\,\,\text{ and }\,\,\,[A,I]=[A,[A,I]]=\operatorname{span}_{\mathbb{C}}\operatorname{FN}_{2}(A,I).$
Proof.
Considering $I$ as a Lie ideal in $A$, it follows from Section 2 that $A[A,I]A$ is the unique two-sided ideal in $A$ to which $I$ is related. Since $I$ is clearly related to itself, we get $I=A[A,I]A$. The inclusion $A[A,I]\subseteq A[A,I]A$ holds since $A$ is unital. The converse inclusion holds since $[A,I]$ is a Lie ideal in $A$; see, for example, [GLT24, Lemma 3.2].
Now the equality $[A,I]=[A,[A,I]]$ also follows from Section 2. Further, since $A$ is zero-product balanced by Lemma 3.3, we have $[A,I]=\operatorname{span}_{\mathbb{C}}\operatorname{FN}_{2}(A,I)$ by Section 3. ∎
[…]
Let $A$ be a unital $C^{*}$-algebra that admits a unital *-homomorphism $M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})\to A$. Applying Section 3 for the ideal $A$, we get
$A=A[A,A]A,$
which means that $A$ is generated by its commutators as an ideal.
In fact, $A$ is even generated by its commutators as a ring. Indeed, by [GT25, Theorem 5.15], every element in $A$ is a sum of a commutator and the product of two commutators, and also the sum of three elements that are products of commutators.
The next result verifies statement (4) in Section 1.
[…]
Let $A$ be a unital $C^{*}$-algebra that admits a unital *-homomorphism $M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})\to A$. There is a natural bijection between the set $\mathcal{I}$ of two-sided ideals in $A$ and the set
:=$\mathcal{L}:=\big{\{}L\subseteq A\text{ Lie ideal}:L=[A,L]\big{\}}$
given by the maps $I\mapsto[A,I]$ and $L\mapsto A[A,L]A$.
Proof.
We denote the maps by $\alpha\colon\mathcal{I}\to\mathcal{L}$ and $\beta\colon\mathcal{L}\to\mathcal{I}$. Given $I\in\mathcal{I}$, using Section 3 at the second step, we have
$\alpha(I)=[A,I]=[A,[A,I]]=[A,\alpha(I)]$
which shows that $\alpha$ is well-defined.
Next, we show that $\alpha$ and $\beta$ are inverses of each other. First, given $I\in\mathcal{I}$, using Section 3 at the last step, we get
$\beta(\alpha(I))=\beta([A,I])=A[A,I]A=I.$
Second, given $L\in\mathcal{L}$, it follows from $L=[A,L]$ that $[A,L]=[A,[A,L]]$, and thus $L=[A,[A,L]]$. Using this at the last step, and using Section 2 at the third step, we get
$\alpha(\beta(L))=\alpha(A[A,L]A)=[A,A[A,L]A]=[A,[A,L]]=L,$
as desired, ∎
Robert showed in [Rob16, Lemma 2.1] that every nilpotent element in a $C^{*}$-algebra is a sum of commutators. With view towards Section 3, we ask:
[…]
Let $I$ be a (not necessarily closed) two-sided ideal in a $C^{*}$-algebra $A$. Does every nilpotent element in $I$ belong to $[A,I]$?
We give positive answers to Section 3 for arbitrary two-sided ideals in von Neumann algebras (Section 6), and for semiprime two-sided ideals in $C^{*}$-algebras (Section 4).
Following [GKT23, Definition 3.2], a Dixmier ideal in a $C^{*}$-algebra $A$ is a two-sided ideal $I\subseteq A$ such that $I$ is positively spanned (that is, $I=\operatorname{span}_{\mathbb{C}}(I\cap A_{+})$) and hereditary (that is, if $a,b\in A_{+}$ satisfy $a\leq b$ and $b\in I$, then $a\in I$) and strongly invariant (that is, if $x\in A$ satisfies $x^{*}x\in I$, then $xx^{*}\in I$). Two-sided ideals in von Neumann algebras are automatically Dixmier ideals, which was essentially shown by Dixmier (see [GKT23, Proposition 3.4]).
Given a Dixmier ideal $I$ and $s\in(0,\infty)$, there is a unique Dixmier ideal whose set of positive elements is $\{a^{s}:a\in I\cap A_{+}\}$, and we use $I^{s}$ to denote this ideal; see Proposition 3.7 and Definition 3.8 in [GKT23]. We note that for $n\in{\mathbb{N}}$ the ring-theoretic definition of $I^{n}$ as the two-sided ideal generated by $x_{1}\cdots x_{n}$ for $x_{1},\ldots,x_{n}\in I$ agrees with the ‘new’ definition of $I^{n}$ as the linear span of elements $a^{n}$ for $a\in I\cap A_{+}$. This leads to a well-behaved theory of roots and powers for Dixmier ideals ([GKT23, Theorem 3.9]), which we recall for the convenience of the reader.
[…]
Let $I\subseteq A$ be a Dixmier ideal in a $C^{*}$-algebra. Then
$(I^{s})^{t}=I^{st}=(I^{t})^{s},\,\,\,\text{ and }\,\,\,I^{s}I^{t}=I^{s+t}=I^{t}I^{s}$
for all $s,t\in(0,\infty)$.
Lemma 3.4.
Let $I\subseteq A$ be a Dixmier ideal in a $C^{*}$-algebra, let $x\in I$ with polar decomposition $x=v|x|$ in $A^{**}$, and let $s\in(0,\infty)$. Then
$v|x|^{s},|x|^{s},|x^{*}|^{s},|x|^{s}v^{*}\in I^{s}.$
Proof.
Since $I$ is a Dixmier ideal, we have $x^{*}\in I$, and thus $x^{*}x,xx^{*}\in I^{2}$. Then by Section 3, we get
$|x|=(x^{*}x)^{\frac{1}{2}}\in(I^{2})^{\frac{1}{2}}=I,\,\,\,\text{ and }\,\,\,|x^{*}|=(xx^{*})^{\frac{1}{2}}\in(I^{2})^{\frac{1}{2}}=I.$
It follows that $|x|^{s},|x^{*}|^{s}\in I^{s}$.
By [GKT23, Proposition 3.10], given a Dixmier ideal $J\subseteq A$, an element $y\in A$ belongs to $J^{\frac{1}{2}}$ if and only if $y^{*}y\in J$. Applying this for the Dixmier ideal $I^{2s}$, and using that
$\big{(}v|x|^{s}\big{)}^{*}\big{(}v|x|^{s}\big{)}=|x|^{2s}\in I^{2s},$
we deduce that $v|x|^{s}\in(I^{2s})^{\frac{1}{2}}=I^{s}$. Since $I^{s}$ is a Dixmier ideal, we further have $|x|^{s}v^{*}=(v|x|^{s})^{*}\in I^{s}$. ∎
The next result showcases a situation when every element in a Dixmier ideal $I$ is a sum of square-zero element in $I$ and products of pairs of square-zero elements in $I^{\frac{1}{2}}$. In Section 6, we strengthen this result for ideals in von Neumann algebras.
[…]
Let $A$ be a unital $C^{*}$-algebra that admits a unital *-homomorphism $M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})\to A$, and let $I\subseteq A$ be a Dixmier ideal. Then
$I=[A,I]+[A,I^{\frac{1}{2}}]^{2}.$
Proof.
We first show that
$A\cdot\operatorname{FN}_{2}(A,I)\subseteq[A,I]+[A,I^{\frac{1}{2}}]^{2}.$
Let $c\in A$ and $x\in\operatorname{FN}_{2}(A,I)$. By Lemma 3.1, we can pick $a,b\in A_{+}$ and $y\in I$ such that $x=ayb$ and $ab=ba=0$. Set $z:=yb\in I$, and let $z=v|z|$ be the polar decomposition in $A^{**}$. Then
$cx=\big{[}ca^{\frac{1}{2}},a^{\frac{1}{2}}z\big{]}+a^{\frac{1}{2}}zca^{\frac{1}{2}}=\big{[}ca^{\frac{1}{2}},a^{\frac{1}{2}}z\big{]}+\big{(}a^{\frac{1}{2}}v|z|^{\frac{1}{2}}\big{)}\big{(}|z|^{\frac{1}{2}}ca^{\frac{1}{2}}\big{)}.$
By Lemma 3.4, $v|z|^{\frac{1}{2}}$ and $|z|^{\frac{1}{2}}$ belong to $I^{\frac{1}{2}}$. Further, since $za=0$, we get $|z|^{\frac{1}{2}}a^{\frac{1}{2}}=0$, and hence
$a^{\frac{1}{2}}v|z|^{\frac{1}{2}}=\big{[}a^{\frac{1}{2}},v|z|^{\frac{1}{2}}\big{]}\in[A,I^{\frac{1}{2}}],\,\,\,\text{ and }\,\,\,|z|^{\frac{1}{2}}ca^{\frac{1}{2}}=\big{[}-ca^{\frac{1}{2}},|z|^{\frac{1}{2}}\big{]}\in[A,I^{\frac{1}{2}}].$
It follows that $cx$ belongs to $[A,I]+[A,I^{\frac{1}{2}}]^{2}$.
Using Section 3 at the first and second step, and using the above at the last step, we get
$I=A[A,I]=\operatorname{span}_{\mathbb{C}}\Big{(}A\cdot\operatorname{FN}_{2}(A,I)\Big{)}\subseteq[A,I]+[A,I^{\frac{1}{2}}]^{2}.$
On the other hand, we clearly have $[A,I]\subseteq I$. Using Section 3, we also have $[A,I^{\frac{1}{2}}]^{2}\subseteq(I^{\frac{1}{2}})^{2}=I$. ∎
4. Square-zero elements in semiprime ideals in C*-algebras
In this section, we study the subspace $N$ generated by the square-zero elements in a semiprime ideal in a $C^{*}$-algebra $A$. We show that $ANA$, the two-sided ideal generated by $N$, satisfies $ANA=[A,N]+N^{2}$; see Section 4. This generalize results from [GT24, Section 4].
We then characterize semiprimeness for two-sided ideals in a $C^{*}$-algebra whose unit is weakly divisible of degree $2$; see Section 4. In particular, such an ideal is semiprime if and only if it is generated by its commutators as an ideal (or as a ring).
We begin with some preparatory results about commutators involving Dixmier ideals, which might be of use in future work.
Lemma 4.1.
Let $I\subseteq A$ be a Dixmier ideal in a $C^{*}$-algebra. Then
$[A,I]\subseteq\big{[}I^{\frac{1}{2}},I^{\frac{1}{2}}\big{]}\subseteq I.$
Proof.
By Section 3, we have $I=I^{\frac{1}{2}}I^{\frac{1}{2}}$. Using that $[a,xy]=[ax,y]+[ya,x]$ for $a\in A$ and $x,y\in I$, it follows that
$[A,I]=\big{[}A,I^{\frac{1}{2}}I^{\frac{1}{2}}\big{]}\subseteq\big{[}AI^{\frac{1}{2}},I^{\frac{1}{2}}\big{]}+\big{[}I^{\frac{1}{2}}A,I^{\frac{1}{2}}\big{]}\subseteq\big{[}I^{\frac{1}{2}},I^{\frac{1}{2}}\big{]}.$
Further, we have $[I^{\frac{1}{2}},I^{\frac{1}{2}}]\subseteq I^{\frac{1}{2}}I^{\frac{1}{2}}=I$. ∎
Recall that $\operatorname{N}_{2}(I)$ denotes the set of square-zero elements in an ideal $I$.
Lemma 4.2.
Let $I\subseteq A$ be a Dixmier ideal in a $C^{*}$-algebra. Then
$\operatorname{N}_{2}(I)\subseteq\operatorname{span}_{\mathbb{C}}\big{[}\operatorname{N}_{2}(I^{\frac{1}{2}}),\operatorname{N}_{2}(I^{\frac{1}{2}})\big{]}.$
Proof.
Let $x\in\operatorname{N}_{2}(I)$, and let $x=v|x|$ be the polar decomposition in $A^{**}$. We have $x=|x^{*}|v=|x^{*}|^{\frac{1}{2}}v|x|^{\frac{1}{2}}$. Since $x^{2}=0$, the elements $|x|$ and $|x^{*}|$ are orthogonal, and hence so are $|x|^{\frac{1}{2}}$ and $|x^{*}|^{\frac{1}{2}}$. Therefore
$x=\big{[}\tfrac{1}{2}v|x|^{\frac{1}{2}},|x|^{\frac{1}{2}}-|x^{*}|^{\frac{1}{2}}\big{]}.$
By Lemma 3.4, the element $\tfrac{1}{2}v|x|^{\frac{1}{2}}$ belongs to $I^{\frac{1}{2}}$, and thus $\tfrac{1}{2}v|x|^{\frac{1}{2}}\in\operatorname{N}_{2}(I^{\frac{1}{2}})$. Set
$y:=|x|^{\frac{1}{2}}-|x^{*}|^{\frac{1}{2}}+v|x|^{\frac{1}{2}}-v^{*}|x^{*}|^{\frac{1}{2}}.$
It is elementary (but tedious) to check that $y^{2}=0$. Note that $y$ belongs to $I^{\frac{1}{2}}$ by Lemma 3.4, and thus $y\in\operatorname{N}_{2}(I^{\frac{1}{2}})$. Using that $v|x|^{\frac{1}{2}}$ and $v^{*}|x^{*}|^{\frac{1}{2}}$ also belong to $N_{2}(I^{\frac{1}{2}})$, it follows that
$|x|^{\frac{1}{2}}-|x^{*}|^{\frac{1}{2}}=y-v|x|^{\frac{1}{2}}+v^{*}|x^{*}|^{\frac{1}{2}}\in\operatorname{span}_{\mathbb{C}}N_{2}(I^{\frac{1}{2}}),$
and thus
$x=\big{[}\tfrac{1}{2}v|x|^{\frac{1}{2}},|x|^{\frac{1}{2}}-|x^{*}|^{\frac{1}{2}}\big{]}\in\operatorname{span}_{\mathbb{C}}\big{[}N_{2}(I^{\frac{1}{2}}),N_{2}(I^{\frac{1}{2}})\big{]},$
as desired. ∎
A two-sided ideal $I$ in a ring $R$ is said to be semiprime if for all two-sided ideals $J\subseteq R$ we have $J\subseteq I$ whenever $J^{2}\subseteq I$. We refer to [Lam01, Section 10] for details. In [GKT23, Section 5], together with Gardella and Kitamura, we showed that a two-sided ideal $I$ in a $C^{*}$-algebra is semiprime if and only if it is idempotent (that is, $I=I^{2}$), if and only if it is a Dixmier ideal and satisfies $I=I^{\frac{1}{2}}$.
The next result is a generalization of [GT24, Theorem 4.2] from the study of square-zero elements in $C^{*}$-algebras to square-zero elements in semiprime ideals in $C^{*}$-algebras. Recall that $\widetilde{A}$ denotes the minimal unitization of a $C^{*}$-algebra $A$.
[…]
Let $I\subseteq A$ be a semiprime ideal in a $C^{*}$-algebra, let $N$ denote the subspace generated by $\operatorname{N}_{2}(I)$, and let $J:=\widetilde{A}N\widetilde{A}$ denote the two-sided ideal of $A$ generated by $N$. Then $J$ is semiprime, and we have
$N=[N,N]\subseteq N^{2},\quad J=[J,J]^{2}=AN=[A,N]^{2}=[A,N]+N^{2},$
and
$[J,J]=[A,J]=[A,[A,N]]=[[J,J],[J,J]]=[[A,N],[A,N]].$
Proof.
We will use that $I$ is a Dixmier ideal satisfying $I=I^{\frac{1}{2}}$. Applying Lemma 4.2 at the second step, we have
$N=\operatorname{span}_{\mathbb{C}}\operatorname{N}_{2}(I)\subseteq\operatorname{span}_{\mathbb{C}}\big{[}\operatorname{N}_{2}(I^{\frac{1}{2}}),\operatorname{N}_{2}(I^{\frac{1}{2}})\big{]}=\operatorname{span}_{\mathbb{C}}\big{[}\operatorname{N}_{2}(I),\operatorname{N}_{2}(I)\big{]}=[N,N].$
The inclusion $[N,N]\subseteq N^{2}$ follows since every commutator in $[N,N]$ belongs to $N^{2}$. To verify that $[N,N]\subseteq N$, let $x,y\in\operatorname{N}_{2}(I)$. Then
$[x,y]=(1+x)y(1-x)-y+xyx.$
Each of the three summands is a square-zero element in $I$, which shows $[x,y]\in N$.
We have
$J:=\widetilde{A}N\widetilde{A}\subseteq\widetilde{A}N^{2}\widetilde{A}\subseteq\big{(}\widetilde{A}N\big{)}\big{(}N\widetilde{A}\big{)}\subseteq J^{2},$
which shows that $J$ is idempotent and therefore semiprime by [GKT23, Theorem A].
The inclusion $AN\subseteq\widetilde{A}N\widetilde{A}=J$ is clear. Conversely, given $a\in\widetilde{A}$, $x\in\operatorname{N}_{2}(I)$ and a unitary $u\in\widetilde{A}$, using that $u^{*}xu\in\operatorname{N}_{2}(I)$ and that $N\subseteq N^{2}$, we have
$axu=au(u^{*}xu)\in\widetilde{A}\operatorname{N}_{2}(I)\subseteq\widetilde{A}N\subseteq\widetilde{A}N^{2}\subseteq AN.$
Since every element in $\widetilde{A}$ is a linear combination of four unitaries, we get $J=AN$.
The inclusion $[A,N]+N^{2}\subseteq\widetilde{A}N\widetilde{A}=J$ is clear. For the converse inclusion, we use a similar argument as in the proof of Section 3 to show that $AN\subseteq[A,N]+N^{2}$. Given $a\in A$ and $x\in\operatorname{N}_{2}(I)$, let $x=v|x|$ be the polar decomposition in $A^{**}$. Since $I$ is semiprime, it is Dixmier and $I=I^{\frac{1}{2}}=I^{\frac{1}{4}}$. By Lemma 3.4, we have $v|x|^{\frac{1}{2}}\in I^{\frac{1}{2}}$ and thus $v|x|^{\frac{1}{2}}\in\operatorname{N}_{2}(I)$. Similarly, we have $v|x|^{\frac{1}{4}},|x|^{\frac{1}{4}}a|x^{*}|^{\frac{1}{2}}\in\operatorname{N}_{2}(I)$. Using that $x=|x^{*}|^{\frac{1}{2}}v|x|^{\frac{1}{2}}$, we get
$ax=\big{[}a|x^{*}|^{\frac{1}{2}},v|x|^{\frac{1}{2}}\big{]}+\big{(}v|x|^{\frac{1}{4}}\big{)}\big{(}|x|^{\frac{1}{4}}a|x^{*}|^{\frac{1}{2}}\big{)}\in[A,\operatorname{N}_{2}(I)]+\operatorname{N}_{2}(I)^{2}.$
Using that $J=AN$, we get $J=[A,N]+N^{2}$.
The inclusion $[A,N]^{2}\subseteq\widetilde{A}N\widetilde{A}=J$ is clear. Conversely, using that $N=[N,N]$, we have
$[A,N]=[A,[N,N]]\subseteq[N,[A,N]]\subseteq[[A,N],[A,N]]\subseteq[A,N]^{2}.$
We also have $N^{2}=[N,N]^{2}\subseteq[A,N]^{2}$ and thus
$J=[A,N]+N^{2}\subseteq[A,N]^{2}.$
The inclusion $[J,J]\subseteq[A,J]$ is clear. Using that $J=J^{2}$ and $[a,xy]=[ax,y]+[ya,x]$ for $a\in A$ and $xy,\in J$, we have
$[A,J]=[A,J^{2}]\subseteq[AJ,J]+[JA,J]\subseteq[J,J].$
This shows that $[J,J]=[A,J]$. We get
$[[A,N],[A,N]]\subseteq[A,[A,N]]\subseteq[A,J]=[J,J].$
We have
$\big{[}[A,N],N^{2}\big{]}\subseteq[A,N^{2}]\subseteq[AN,N]+[NA,N]\subseteq[A,N].$
and similarly $[N^{2},N^{2}]\subseteq[A,N^{2}]\subseteq[A,N]$. We also have
$[A,N]=[A,[N,N]]\subseteq[N,[A,N]]=[[N,N],[A,N]]\subseteq[[A,N],[A,N]].$
Using that $J=[A,N]+N^{2}$, it follows that
$[J,J]$ $\subseteq\big{[}[A,N],[A,N]\big{]}+\big{[}N^{2},[A,N]\big{]}+\big{[}[A,N],N^{2}\big{]}+\big{[}N^{2},N^{2}\big{]}$
$\subseteq\big{[}[A,N],[A,N]\big{]}.$
Using that $[A,N]\subseteq[A,J]=[J,J]$, we also have
$[J,J]=\big{[}[A,N],[A,N]\big{]}\subseteq\big{[}[J,J],[J,J]\big{]}.$
Since the converse inclusion is clear.
Finally, since $[A,N]\subseteq[J,J]$ and $J=[A,N]^{2}$, we have $J=[J,J]^{2}$. ∎
Lemma 4.3.
Let $I\subseteq A$ be a Dixmier ideal in a $C^{*}$-algebra, and let $x\in I$ be nilpotent. Then $x\in[I^{\frac{1}{2}},I^{\frac{1}{2}}]$.
Proof.
As in the proof of [Rob16, Lemma 2.1], we proceed by induction on the degree of nilpotency. To start, let $x\in I$ be a square-zero element. By Lemma 3.4, the elements $v|x|^{\frac{1}{2}}$ and $|x|^{\frac{1}{2}}$ belong to $I^{\frac{1}{2}}$, and thus
$x=\big{[}v|x|^{\frac{1}{2}},x^{\frac{1}{2}}\big{]}\in\big{[}I^{\frac{1}{2}},I^{\frac{1}{2}}\big{]}.$
Next, assume that for some $k\geq 2$ we have shown that every element $x\in I$ with $x^{k}=0$ belongs to $[I^{\frac{1}{2}},I^{\frac{1}{2}}]$. Let $x\in I$ satisfy $x^{k+1}=0$, and let $x=v|x|$ be the polar decomposition of $x$ in $A^{**}$. Set $y:=|x|^{\frac{1}{2}}v|x|^{\frac{1}{2}}$, which belongs to $A$. As in the proof of [Rob16, Lemma 2.1], one shows that $y^{k}=0$. By Lemma 3.4, we have $v|x|^{\frac{1}{2}},|x|^{\frac{1}{2}}\in I^{\frac{1}{2}}$. By assumption of the induction, we have $y\in[I^{\frac{1}{2}},I^{\frac{1}{2}}]$, and thus
$x=\big{[}v|x|^{\frac{1}{2}},|x|^{\frac{1}{2}}\big{]}+y\in\big{[}I^{\frac{1}{2}},I^{\frac{1}{2}}\big{]}$
as desired. ∎
[…]
Let $I\subseteq A$ be a semiprime ideal in a $C^{*}$-algebra, and let $x\in I$ be nilpotent. Then $x\in[A,I]$.
Proof.
Since $I$ is semiprime, it is a Dixmier ideal and satisfies $I=I^{\frac{1}{2}}$. Applying Lemma 4.3, we have
$x\in[I^{\frac{1}{2}},I^{\frac{1}{2}}]\subseteq[A,I^{\frac{1}{2}}]=[A,I],$
as desired. ∎
[…]
Let $A$ be a unital $C^{*}$-algebra that admits a unital *-homomorphism $M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})\to A$, and let $I\subseteq A$ be a two-sided ideal. The following statements are equivalent:
  1. (1)
    The ideal $I\subseteq A$ is semiprime.
  2. (2)
    We have $I=A[I,I]A$, that is, $I$ is generated by its commutators as an ideal in $A$.
  3. (3)
    We have $I=\widetilde{I}[I,I]\widetilde{I}$, that is, $I$ is generated by its commutators as an ideal.
  4. (4)
    We have $I=[I,I]^{2}$.
  5. (5)
    We have $[I,I]=[[I,I],[I,I]]$.
Proof.
To show that (2) implies (1), assume that $I$ is generated by $[I,I]$ as an ideal in $A$. We clearly have $[I,I]\subseteq I^{2}$, and therefore
$I=A[I,I]A\subseteq AI^{2}A=I^{2}.$
It follows that $I$ is idempotent, and thus semiprime by [GKT23, Theorem A].
It is clear that (4) implies (3), which in turn implies (2).
Assume that $I$ is semiprime, and let $N$ denote the subspace generated by the square-zero elements in $I$. Applying Section 3, we have
$I=A[A,I]A=\operatorname{span}_{\mathbb{C}}A\operatorname{FN}_{2}(A,I)A\subseteq ANA.$
Now it follows from Section 4 that
$I=ANA=[ANA,ANA]^{2}=[I,I]^{2},$
which verifies (4), and that
$[I,I]=[ANA,ANA]=[[ANA,ANA],[ANA,ANA]]=[[I,I],[I,I]],\,\,\,\text{ and }\,\,\,$
which verifies (5).
Finally, assuming (5) let us verify (1). Using that $[a,xy]=[ax,y]+[ya,x]$ for $a\in A$ and $x,y\in I$, we have
$[A,I^{2}]\subseteq[AI,I]+[IA,I]\subseteq[I,I].$
Applying Section 3 for the ideal $I^{2}$, we get
$I^{2}=A[A,I^{2}]A\subseteq A[I,I]A\subseteq A[[I,I],[I,I]]A\subseteq I^{4}.$
It follows that $I^{2}=I^{4}$, and thus semiprime by [GKT23, Theorem A]. ∎
5. Lie ideals in properly infinite C*-algebras
Let $A$ be a $C^{*}$-algebra with a unit that is weakly divisible of degree $2$, that is, that admits a unital *-homomorphism $M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})\to A$, and let $L\subseteq A$ be a (not necessarily closed) Lie ideal. We have seen in Section 2 that $L$ is related to the two-sided ideal $I:=A[A,L]A$. Further, by Section 2, $L$ is commutator equivalent to $I$ if and only if $[A,L]=[A,[A,L]]$.
This raises the question when Lie ideals $L$ in $A$ satisfy $[A,L]=[A,[A,L]]$. A sufficient condition is that $A=Z(A)+[A,A]$, where $Z(A)$ denotes the center of $A$; see Lemma 5.1. We observe in Section 5 that this condition is satisfies in a number of situations, in particular for all properly infinite $C^{*}$-algebras.
We then prove the main result of the paper: Every Lie ideal in a unital, properly infinite $C^{*}$-algebra is commutator equivalent to a (unique) two-sided ideal; see Section 5.
Lemma 5.1.
Let $A$ be a unital $C^{*}$-algebra such that $A=Z(A)+[A,A]$, and let $L\subseteq A$ be a Lie ideal. Then
$[A,L]=[A,[A,L]].$
Proof.
The inclusion $\supseteq$’ is clear. Conversely, using the assumption at the first step, and using the Jacobi identity at the last step, we get
$[A,L]=\big{[}Z(A)+[A,A],L\big{]}=\big{[}[A,A],L\big{]}\subseteq\big{[}A,[A,L]\big{]},$
as desired. ∎
Following Winter [Win12], a $C^{*}$-algebra $A$ is said to be pure if its Cuntz semigroup $\operatorname{Cu}(A)$ is almost unperforated and almost divisible. Pure $C^{*}$-algebras form a robust class [APTV24] that includes every $C^{*}$-algebra that tensorially absorbs the Jiang-Su algebra $\mathcal{Z}$, that is, $A\cong A\otimes\mathcal{Z}$. By [APT18, Theorem 7.3.11], a $C^{*}$-algebra $A$ is pure if and only if $\operatorname{Cu}(A)\cong\operatorname{Cu}(A)\otimes\operatorname{Cu}(\mathcal{Z})$.
By a quasitrace of a $C^{*}$-algebra $A$ we mean a lower-semicontinuous, $[0,\infty]$-valued $2$-quasitrace. By Haagerup’s theorem, every bounded quasitrace on a unital, exact $C^{*}$-algebra is a trace, and this was later extended by Kirchberg, who showed that arbitrary quasitraces on exact $C^{*}$-algebras are traces.
A unital $C^{*}$-algebra $A$ is weakly central if the center separates maximal ideals: Maximal ideals $M_{1},M_{2}\subseteq A$ satisfy $M_{1}=M_{2}$ whenever $M_{1}\cap Z(A)=M_{2}\cap Z(A)$. By Vesterstrøm’s theorem, $A$ is weakly central if and only if $A$ has the centre-quotient property: For every closed, two-sided ideal $I\subseteq A$, the quotient map $A\to A/I$ maps $Z(A)$ onto $Z(A/I)$.
A unital $C^{*}$-algebra $A$ has the Dixmier property if for every element $a\in A$ the closed convex hull of the unitary orbit $\{uau^{*}:u\in A\text{ unitary}\}$ has nonempty intersection with the center $Z(A)$. It was shown in [ART17, Theorem 1.1] that this property holds if and only if $A$ is weakly central, every simple quotient of $A$ has at most one tracial state, and every extreme tracial state of $A$ factors through a simple quotient. Dixmier showed that every von Neumann algebra has the property named after him.
[…]
Let $A$ be a unital $C^{*}$-algebra. Then $A=Z(A)+[A,A]$ in each of the following cases:
  1. (1)
    If $A$ is pure, every quasitrace on $A$ is a trace, and $A$ has the Dixmier property.
  2. (2)
    If $A$ has no tracial states (in particular, if $A$ is properly infinite).
  3. (3)
    If $A$ is a von Neumann algebra.
Proof.
(1) If $A$ has the Dixmier property, then every element in $A$ is a Dixmier element in the sense of [AGR23], and thus it follows from [AGR23, Lemma 4.10] that $A=Z(A)+\overline{[A,A]}$.
If $A$ is pure and every quasitrace on $A$ is a trace, then $[A,A]$ is closed by [NR16, Theorem 1.1]. Combining both results, we get $A=Z(A)+[A,A]$.
(2) By [Pop02, Theorem 1], we have $A=[A,A]$ if (and only if) $A$ has no tracial states.
(3) It was shown in the proof of [BKS08, Theorem 5.19] that every von Neumann algebra $A$ satisfies $A=Z(A)+[A,A]$. ∎
[…]
Let $A$ be a unital, properly infinite $C^{*}$-algebra, and let $L\subseteq A$ be a Lie ideal. Then for the two-sided ideal $I:=A[A,L]A$, we have
$[A,L]=[A,[A,L]]=[A,I].$
In particular, $L$ is commutator equivalent to $I$, and hence also embraced by $I$, and related to $I$.
Further, $I$ is the only two-sided ideal of $A$ to which $L$ is related.
Proof.
Since $A$ is unital and properly infinite, there exists a unital *-homomorphism from the Cuntz algebra $\mathcal{O}_{\infty}$ to $A$; see [Bla06, Proposition III.1.3.3]. Further, it is well-known that $\mathcal{O}_{\infty}$ admits a unital *-homomorphism from $M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})$, using for example Section 2 and that $\mathcal{O}_{\infty}$ is a unital $C^{*}$-algebra of real rank zero that admits no characters. It follows that $A$ admits a unital *-homomorphism $M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})\to A$.
We may therefore apply Section 2 and obtain that $L$ is related to $I$, and that $I$ is the only two-sided ideal to which $L$ is related. Further, since $A$ is unital and properly infinite, we have $A=[A,A]$ by Pop’s theorem [Pop02], and thus $[A,L]=[A,[A,L]]$ by Lemma 5.1. It therefore follows from Section 2 that $L$ is commutator equivalent to $I$ (and thus also embraced by $I$). ∎
[…]
Let $A$ be a unital, properly infinite $C^{*}$-algebra. Then an additive subgroup $L\subseteq A$ is a Lie ideal if and only if $[A,I]\subseteq L\subseteq T([A,I])$ for some (unique) two-sided ideal $I\subseteq A$. We get a decomposition of the family $\mathcal{L}$ of all Lie ideals in $A$ as:
$\mathcal{L}=\bigsqcup_{I\in\mathcal{I}}\big{\{}L\subseteq A\text{ additive subgroup}:[A,I]\subseteq L\subseteq T([A,I])\big{\}}$
where $\mathcal{I}$ denotes the lattice of two-sided ideals in $A$.
[…]
The proof of Section 5 can easily be adapted to show the following: In a $C^{*}$-algebra $A$ that admits a unital *-homomorphism $M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})\to A$ and such that $A=Z(A)+[A,A]$, every Lie ideal $L$ is commutator equivalent to the two-sided ideal $A]A,L]A$.
We note that a unital $C^{*}$-algebra admits no characters whenever it is pure or has no tracial stetes. Therefore, with view towards Section 2 and Section 5, we expect that every Lie ideal $L$ is commutator equivalent to $A[A,L]A$ whenever $A$ is a unital, pure, $C^{*}$-algebra with the Dixmier property and such that every quasitrace on $A$ is a trace, or whenever $A$ is unital and has no tracial states.
[…]
Given Lie ideals $L,M\subseteq A$ in a unital, properly infinite $C^{*}$-algebra, what is the relationship between the two-sided ideals generated by $[L,M]$, by $[A,[L,M]]$, and by $[A,L]$ and $[A,M]$?
6. Lie ideals and two-sided ideals in von Neumann algebras
In this section, we show that every (not necessarily closed) Lie ideal $L$ in a von Neumann algebra $M$ is commutator equivalent to the two-sided ideal $I:=M[M,L]M$, that is, $[M,L]=[M,I]$; see Section 6. This recovers [BKS08, Theorem 5.19]. Moreover, if $M$ has zero commutative summand, then we show that $I$ is the only two-sided ideal to which $L$ is related, which solves [BKS08, Problem 5.21].
Given a two-sided ideal $I$ in a von Neumann algebra with zero commutative summand, every element in $I$ is a sum of products of pairs of square-zero elements in $I^{\frac{1}{2}}$; see Section 6.
[…]
Let $M$ be a von Neumann algebra, and let $L\subseteq M$ be a Lie ideal. Then, for the two-sided ideal $I:=M[M,L]M$, we have
$[M,L]=[M,[M,L]]=[M,I].$
In particular, $L$ is commutator equivalent to $I$, and hence also embraced by $I$, and related to $I$.
If $M$ has zero commutative summand, then $I$ is the only two-sided ideal of $M$ to which $L$ is related.
Proof.
Let $M=M_{0}\oplus M_{1}$ be the (unique) decomposition of $M$ into a commutative summand $M_{0}$ and a summand $M_{1}$ that admits no characters. By Section 2, there is a unital *-homomorphism $M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})\to M_{1}$. Set
$L_{1}:=M_{1}\cap L,\,\,\,\text{ and }\,\,\,I_{1}:=M_{1}\cap I.$
Then $L_{1}$ is a Lie ideal in $M_{1}$, and we have
$[M,L]=[M_{1},L_{1}],\,\,\,\text{ and }\,\,\,I=I_{1}=M_{1}[M_{1},L_{1}]M_{1}.$
We may apply Section 2 for the Lie ideal $L_{1}$ in $M_{1}$ and obtain that $L_{1}$ is related to $I_{1}$, and that $I_{1}$ is the only two-sided ideal of $M_{1}$ to which $L_{1}$ is related. It follows that $L$ is related to $I$, and if $M_{0}$ is zero, then $I$ is also unique with this property. (If $M_{0}$ is nonzero, then $M_{0}\oplus I$ is another two-sided ideal of $M$ to which $L$ is related.)
As shown in the proof of [BKS08, Theorem 5.19], we have $M=Z(M)+[M,M]$. It follows that $[M,L]=[M,[M,L]]$ by Lemma 5.1, and then
$[M,L]=[M,[M,L]]=[M,I].$
by Section 2. ∎
Recall that $\operatorname{N}_{2}(I)$ denotes the square-zero elements, and we use $\operatorname{Nil}(I)$ to denote the set of nilpotent elements in an ideal $I$. Recall the definition of $\operatorname{FN}_{2}(M,I)$ from Section 3.
[…]
Let $I\subseteq M$ be a two-sided ideal in a von Neumann algebra. Then
$[M,I]=\operatorname{span}_{\mathbb{C}}\operatorname{N}_{2}(I)=\operatorname{span}_{\mathbb{C}}\operatorname{Nil}(I)=\operatorname{span}_{\mathbb{C}}\operatorname{FN}_{2}(M,I).$
Proof.
Let $M=M_{0}\oplus M_{1}$ denote the (unique) decomposition of $M$ into a commutative summand $M_{0}$ and a summand $M_{1}$ that admits no characters. Then
$[M,I]$ $=[M_{1},M_{1}\cap I],\quad\operatorname{N}_{2}(I)=\operatorname{N}_{2}(M_{1}\cap I),$
$\operatorname{Nil}(I)$ $=\operatorname{Nil}(M_{1}\cap I),\,\,\,\text{ and }\,\,\,\operatorname{FN}_{2}(M,I)=\operatorname{FN}_{2}(M_{1},M_{1}\cap I),$
which allows us to reduce the problem to studying the ideal $M_{1}\cap I$ in $M_{1}$.
Without loss of generality, we may therefore assume that $M$ has zero commutative summand. Then $M$ admits a unital *-homomorphism $M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})\to M$ by Section 2. Using Section 3 at the first step, we get
$[M,I]=\operatorname{span}_{\mathbb{C}}\operatorname{FN}_{2}(M,I)\subseteq\operatorname{span}_{\mathbb{C}}\operatorname{N}_{2}(I)\subseteq\operatorname{span}_{\mathbb{C}}\operatorname{Nil}(I).$
It remains to show that $\operatorname{Nil}(I)\subseteq[M,I]$. As in the proof of Lemma 4.3 (and [Rob16, Lemma 2.1]), we proceed by induction on the degree of nilpotency. To start, let $x\in I$ be a square-zero element. Let $x=v|x|$ be the polar decomposition in $M$. Then $|x|=v^{*}x\in I$. Since $x^{2}=0$, we have $|x|v=0$, and thus
$x=\big{[}v,|x|\big{]}\in[M,I].$
Next, assume that for some $k\geq 2$ we have shown that every element $x\in I$ with $x^{k}=0$ belongs to $[M,I]$. Let $x\in I$ satisfy $x^{k+1}=0$. Again, let $x=v|x|$ be the polar decomposition in $M$. We will show that the element $y:=|x|v$ satisfies $y^{k}=0$. Since $x^{k+1}=0$ and $|x|=v^{*}v|x|$, we have
$y^{k}|x|=|x|v\cdots|x|v|x|=v^{*}v|x|v\cdots|x|v|x|=v^{*}x^{k+1}=0.$
For every polynomial $p$ with vanishing constant term it follows that $y^{k}p(|x|)=0$. Using functional calculus, we get $y^{k}|x|^{\frac{1}{n}}$ for every $n\geq 1$. The sequence $(|x|^{\frac{1}{n}})_{n}$ converges in the weak*-topology of $M$ to $v^{*}v$, which is the support projection of $|x|$. We deduce that $y^{k}v^{*}v=0$. Using that $v=vv^{*}v$, we get
$y^{k}=y^{k-1}|x|v=y^{k-1}|x|vv^{*}v=y^{k}v^{*}v=0.$
By assumption of the induction, we have $y\in[M,I]$, and thus
$x=\big{[}v,|x|\big{]}+|x|v\in[M,I],$
as desired. ∎
[…]
Let $H$ be a separable, infinite-dimensional Hilbert space, and let $I$ be a two-sided ideal in the von Neumann algebra $\mathcal{B}(H)$. Given a nilpotent element $x\in I$, it follows from Section 6 that $x\in[\mathcal{B}(H),I]$, that is, $x$ is a finite sum of commutators of an element in $\mathcal{B}(H)$ and an element in $I$. Dykema and Krishnaswamy-Usha showed in [DKU18, Proposition 3.1] that one summand suffices, that is, we have $x=[y,z]$ for some $y\in\mathcal{B}(H)$ and $z\in I$. Does their result hold for nilpotent elements in two-sided ideals in arbitrary von Neumann algebras?
[…]
Let $M$ be a von Neumann algebra with zero commutative summand, and let $I\subseteq M$ be a two-sided ideal. Then
$I=\big{[}M,I^{\frac{1}{2}}\big{]}^{2}=\big{[}I^{\frac{1}{4}},I^{\frac{1}{4}}\big{]}^{2},$
and
$[M,I]\subseteq\big{[}[M,I^{\frac{1}{2}}],[M,I^{\frac{1}{2}}]\big{]}\subseteq\big{[}[I^{\frac{1}{4}},I^{\frac{1}{4}}],[I^{\frac{1}{4}},I^{\frac{1}{4}}]\big{]}\subseteq[I^{\frac{1}{2}},I^{\frac{1}{2}}].$
Proof.
We will use that $I$ is a Dixmier ideal, since two-sided ideals in von Neumann algebras are automatically Dixmier ideals ([GKT23, Proposition 3.4]). Using Section 6 at the first and last step, and using Lemma 4.2 at the second step, we have
$[M,I]=\operatorname{span}_{\mathbb{C}}\operatorname{N}_{2}(I)\subseteq\operatorname{span}_{\mathbb{C}}\big{[}\operatorname{N}_{2}(I^{\frac{1}{2}}),\operatorname{N}_{2}(I^{\frac{1}{2}})\big{]}=\big{[}[M,I^{\frac{1}{2}}],[M,I^{\frac{1}{2}}]\big{]}.$
Using Lemma 4.1, we then get
$[M,I]\subseteq\big{[}[M,I^{\frac{1}{2}}],[M,I^{\frac{1}{2}}]\big{]}\subseteq\big{[}[I^{\frac{1}{4}},I^{\frac{1}{4}}],[I^{\frac{1}{4}},I^{\frac{1}{4}}]\big{]}\subseteq\big{[}I^{\frac{1}{2}},I^{\frac{1}{2}}\big{]}$
By Section 2, there is a unital *-homomorphism $M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})\to M$. We can therefore apply Section 3 at the first step, and together with the above we obtain that
$I=[M,I]+\big{[}M,I^{\frac{1}{2}}\big{]}^{2}\subseteq\big{[}M,I^{\frac{1}{2}}\big{]}^{2}.$
Applying Lemma 4.1 for the Dixmier ideal $I^{\frac{1}{2}}$, and using Section 3 at the last step, we get
$I\subseteq\big{[}M,I^{\frac{1}{2}}\big{]}^{2}\subseteq\big{[}I^{\frac{1}{4}},I^{\frac{1}{4}}\big{]}^{2}\subseteq\big{(}I^{\frac{1}{2}}\big{)}^{2}=I,$
as desired. ∎
[…]
It is likely that the inclusions involving $[M,I]$ in Section 6 are equalities. The key question is whether $[I^{\frac{1}{2}},I^{\frac{1}{2}}]$ is a subset of $[M,I]$. Using that $I$ is a Dixmier ideal and therefore admits a well-behaved theory of roots and powers (Section 3), we have
$\big{[}I^{\frac{1}{2}},I^{\frac{1}{2}}\big{]}=\big{[}I^{\frac{1}{4}}I^{\frac{1}{4}},I^{\frac{1}{2}}\big{]}\subseteq\big{[}I^{\frac{1}{4}},I^{\frac{1}{4}}I^{\frac{1}{2}}\big{]}+\big{[}I^{\frac{1}{2}},I^{\frac{1}{2}}I^{\frac{1}{4}}\big{]}=\big{[}I^{\frac{1}{4}},I^{\frac{3}{4}}\big{]},$
and by an analogous argument $[I^{\frac{1}{4}},I^{\frac{3}{4}}]\subseteq[I^{\frac{1}{8}},I^{\frac{7}{8}}]$. Inductively, it follows that $[I^{\frac{1}{2}},I^{\frac{1}{2}}]\subseteq[I^{\frac{1}{2^{n}}},I^{\frac{2^{n}-1}{2^{n}}}]$ for every $n\geq 1$, and thus
$\big{[}I^{\frac{1}{2}},I^{\frac{1}{2}}\big{]}\subseteq\bigcap_{n\geq 1}\big{[}I^{\frac{1}{2^{n}}},I^{\frac{2^{n}-1}{2^{n}}}\big{]}\subseteq\bigcap_{\varepsilon>0}\big{[}M,I^{1-\varepsilon}\big{]}.$
It is, however, not clear if $\bigcap_{\varepsilon>0}[M,I^{1-\varepsilon}]=[M,I]$.
For two-sided ideals $I,J\subseteq\mathcal{B}(H)$, it was shown in [DFWW04, Theorem 5.10] that $[\mathcal{B}(H),IJ]=[I,J]$. In particular, we have $[\mathcal{B}(H),I]=[I^{\frac{1}{2}},I^{\frac{1}{2}}]$, and we obtain the following consequence of Section 6:
[…]
Let $H$ be a separable, infinite-dimensional Hilbert space, and let $I$ be a two-sided ideal in $\mathcal{B}(H)$. Then
$[\mathcal{B}(H),I]=\big{[}[\mathcal{B}(H),I^{\frac{1}{2}}],[\mathcal{B}(H),I^{\frac{1}{2}}]\big{]}=\big{[}[I^{\frac{1}{4}},I^{\frac{1}{4}}],[I^{\frac{1}{4}},I^{\frac{1}{4}}]\big{]}=\big{[}I^{\frac{1}{2}},I^{\frac{1}{2}}\big{]}.$
[…]
Let $I,J\subseteq M$ be two-sided ideals in a von Neumann algebra. Do we have $[M,IJ]=[I,J]$?