### v2

# Lie ideals in properly infinite C*-algebras

Abstract.
We show that every Lie ideal in a unital, properly infinite $C^{*}$-algebra is commutator equivalent to a unique two-sided ideal. It follows that the Lie ideal structure of such a $C^{*}$-algebra is concisely encoded by its lattice of two-sided ideals. This answers a question of Robert in this setting.
We obtain similar structure results for Lie ideals in unital, real rank zero $C^{*}$-algebras without characters. As an application, we show that every Lie ideal in a von Neumann algebra is related to a unique two-sided ideal, which solves a problem of Brešar, Kissin, and Shulman.
Key words and phrases:
Lie ideals, $C^{*}$-algebras, properly infinite, von Neumann algebras, commutators, square-zero elements
2020 Mathematics Subject Classification:
Primary 46L05, 46L10. Secondary 16W10, 17B60, 47B47.
The first named author was partially supported by the Knut and Alice Wallenberg Foundation (KAW 2021.0140).
1. Introduction
A Lie ideal in a $C^{*}$-algebra $A$ is a linear subspace $L\subseteq A$ such that for every $a\in A$ and $x\in L$ the commutator $[a,x]:=ax-xa$ belongs to $L$. The study of Lie ideals in $C^{*}$-algebras has a long history, and we refer to Section 2 in [Mar10] for an overview of available results and directions of active research.
Immediate examples of Lie ideals in a $C^{*}$-algebra $A$ are all subspaces contained in the center $Z(A)$ and all subspaces containing the commutator subspace $[A,A]$. If $A$ is unital and simple, then this does in fact describes all Lie ideals – a consequence of Herstein’s description of Lie ideals in simple rings [Her55]. Since the center of a unital, simple $C^{*}$-algebra $A$ contains only scalar multiples of the identity, a subspace $L$ is a Lie ideal in $A$ if and only if $L=\{0\}$, $L={\mathbb{C}}1$, or $[A,A]\subseteq L$.
Pop [Pop02] showed that a unital $C^{*}$-algebra satisfies $A=[A,A]$ if and only if $A$ has no tracial states. Consequently, a simple, unital $C^{*}$-algebra $A$ without tracial states has exactly three Lie ideals: $\{0\}$, ${\mathbb{C}}1$ and $A$. Cuntz and Pedersen [CP79] gave a description of the closure $\overline{[A,A]}$ as the intersection of the kernels of all tracial states on $A$. In particular, a simple, unital $C^{*}$-algebra $A$ with a unique tracial state has exactly four closed Lie ideals: $\{0\}$, ${\mathbb{C}}1$, $\overline{[A,A]}$ and $A$. This was first observed by Marcoux and Murphy [MM98], generalizing the case of UHF-algebras, which was first proved by Marcoux [Mar95].
Ng and Robert [NR16] showed that $[A,A]$ is closed in every pure, exact $C^{*}$-algebra. It follows that a simple, unital, exact, pure $C^{*}$-algebra $A$ with a unique tracial state has exactly four Lie ideals: $\{0\}$, ${\mathbb{C}}1$, $[A,A]$ and $A$. This applies for example to UHF-algebras, to irrational rotation algebras, and to the Jiang-Su algebra (where it was first shown by Ng [Ng12]). It also applies to many, and possibly all, simple, reduced group $C^{*}$-algebras. Indeed, if $G$ is a countable, exact, acylindrically hyperbolic group with trivial finite radical and with the rapid decay property, then $C^{*}_{\mathrm{red}}(G)$ has strict comparison [AGKEP24], and therefore is pure [APTV24], and has a unique tracial state [BKKO17].
Turning towards Lie ideals in non-simple $C^{*}$-algebras, it is clear that the structure of the lattice of two-sided ideals will play a role. A systematic study of the interplay between Lie and two-sided ideals was initiated by Brešar, Kissin, and Shulman [BKS08] We first introduce the necessary notation. Following the usual convention, given subspaces $M,N\subseteq A$, we use $[M,N]$ to denote the subspace of $A$ generated by the set of commutators $[x,y]$ for $x\in M$ and $y\in N$. Further, given a subspace $K\subseteq A$, we set
$T(K):=\big{\{}a\in A:[A,a]\subseteq K\big{\}}.$
This ‘derived subspace’ was first considered by Herstein [Her55], and it is denoted by $N(K)$ in [BKS08].
Given a two-sided ideal $I\subseteq A$, Brešar, Kissin, and Shulman observed that a subspace $L\subseteq A$ is a Lie ideal whenever
$[A,I]\subseteq L\subseteq T([A,I]).$
We say that a Lie ideal $L$ is embraced by $I$ if it satisfies these inclusions. The Lie ideals that are embraced by some two-sided ideal are thought of as ‘unsurprising’. A major question is if there exist any Lie ideals that do not arise this way:
[…]
Given a $C^{*}$-algebra $A$, is every Lie ideal of $A$ embraced by some two-sided ideal of $A$?
To study this problem, two further notions were introduced in [BKS08]: Given a Lie ideal $L\subseteq A$ and a two-sided ideal $I\subseteq A$, we say that $L$ is commutator equivalent to $I$ if
$[A,I]=[A,L].$
Further, we say that $L$ is related to $I$ if
$[A,I]\subseteq L\subseteq T(I).$
It is easy to see that these constitute a sufficient and a necessary conditions for embracement. Indeed, if $L$ is commutator equivalent to $I$, then $L$ is embraced by $I$. Further, if $L$ is embraced by $I$, then $L$ is related to $I$.
It was shown in [BKS08, Theorem 5.27] that every Lie ideal $L$ in a $C^{*}$-algebra $A$ is topologically commutator equivalent to a two-sided ideal $I$ in the sense that $[A,L]$ and $[A,I]$ have the same closure. This leads to a description of all closed Lie ideals in terms of closed two-sided ideals. However, this does not show that all closed Lie ideals in a $C^{*}$-algebra are commutator equivalent to (let alone embraced by or related to) a two-sided ideal. For general non-closed Lie ideals even less is known, and Section 1 as well as the following stronger problem remain open:
[…]
Given a $C^{*}$-algebra $A$, is every Lie ideal of $A$ commutator equivalent to some two-sided ideal of $A$?
The results mentioned at the beginning of the introduction show that every Lie ideal in a unital, simple $C^{*}$-algebra $A$ is commutator equivalent to a two-sided ideal, which gives positive solutions to Problems 1 and 1 in this setting. Indeed, the only two-sided ideals are $\{0\}$ and $A$, and a Lie ideal is commutator equivalent to $\{0\}$ if and only if it is contained in $Z(A)$, while a Lie ideal it is commutator equivalent to $A$ if and only if it contains $[A,A]$. (This uses that $[A,A]=[A,[A,A]]$, which follows for example from [GT24, Corollary 3.4].)
In [FMS82], Fong, Miers and Sourour showed that every Lie ideal in the von Neumann algebra of bounded, linear operators on a separable Hilbert space is embraced by a two-sided ideal. This was substantially extended by Brešar, Kissin, and Shulman, who showed in [BKS08, Theorem 5.19] that every Lie ideal in a von Neumann algebra is commutator equivalent to a two-sided ideal, thereby providing positive solutions to Problems 1 and 1 for von Neumann algebras.
Our main result is a positive solution to Section 1 for unital, properly infinite $C^{*}$-algebras.
[…]
Let $L$ be a Lie ideal in a unital, properly infinite $C^{*}$-algebra $A$. Then $L$ is commutator equivalent to the two-sided ideal $I:=A[A,L]A$, and hence also embraced by $I$, and related to $I$. Further, $I$ is the only two-sided ideal of $A$ to which $L$ is related.
The proof of Section 1 is inspired by a result of Marcoux [Mar95], who showed that if $B$ is a unital $C^{*}$-algebra and $n\geq 2$, then every Lie ideal in the matrix algebra $A=M_{n}(B)$ is related to a two-sided ideal in $A$. Our main technical advancement is a generalization of this result to unital $C^{*}$-algebras $A$ that admit a unital *-homomorphism $M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})\to A$.
The next result summarizes Theorems 2, 2, 3, and 3
[…]
Let $A$ be a unital $C^{*}$-algebra that admits a unital *-homomorphism $M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})\to A$. The following statements hold:
  1. (1)
    Given a Lie ideal $L\subseteq A$, for the two-sided ideal $I=A[A,L]A$, we have
    $[A,I]=[A,[A,L]]\subseteq[A,L]\subseteq I,\,\,\,\text{ and }\,\,\,[A,I]\subseteq L\subseteq T(I).$
    Further, $I$ is the only two-sided ideal of $A$ to which $L$ is related.
  2. (2)
    A Lie ideal $L\subseteq A$ is embraced by some two-sided ideal (which then necessarily is $A[A,L]A$) if and only if $L$ is commutator equivalent to some two-sided ideal, if and only if $[A,L]=[A,[A,L]]$.
  3. (3)
    Given a two-sided ideal $I\subseteq A$, we have
    $I=A[A,I]A,\,\,\,\text{ and }\,\,\,[A,I]=[A,[A,I]].$
  4. (4)
    There is a natural bijection between the set $\mathcal{I}$ of two-sided ideals in $A$ and the set
    =$\mathcal{L}=\big{\{}L\subseteq A\text{ Lie ideal}:L=[A,L]\big{\}}.$
    given by the maps $I\mapsto[A,I]$ and $L\mapsto A[A,L]A$.
We note that Section 1 applies to all unital, properly infinite $C^{*}$-algebras, which then yields Section 1. It also applies to unital, real rank zero $C^{*}$-algebras without characters, and in particular to all von Neumann algebras with zero commutative summand, which we use to recover the solution to Section 1 for von Neumann algebras from [BKS08]. Our approach also shows that a Lie ideal in a von Neumann algebra without commutative summand is related to a unique two-sided ideal, which solves [BKS08, Problem 5.21].
[…]
Let $L$ be a Lie ideal in a von Neumann algebra $M$. Then $L$ is commutator equivalent to the two-sided ideal $I:=M[M,L]M$, and hence also embraced by $I$, and related to $I$.
If $M$ has zero commutative summand, then $I$ is the only two-sided ideal of $M$ to which $L$ is related.
Conventions
Given a $C^{*}$-algebra, we use $A_{+}$ to denote its collection of positive elements, and we use $\widetilde{A}$ to denote the minimal unitization. The absolute value of an element $x\in A$ is defined as $|x|:=(x^{*}x)^{\frac{1}{2}}$.
Acknowledgements
The author thanks Leonel Robert for valuable comments regarding the Dixmier property for $C^{*}$-algebras, and Matej Brešar for helpful feedback on an earlier version of this paper. The author is also grateful to the anonymous referees for their insightful suggestions, in particular for pointing out a consequence of the results that is now included as Section 2.
2. Lie ideals in C*-algebras with weakly divisible unit
In this section, we prove that every Lie ideal $L$ in a $C^{*}$-algebra $A$ that admits a unital *-homomorphism $M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})\to A$ is related to a unique two-sided ideal $I$ in $A$; see Section 2. We further show that $L$ is commutator equivalent to $I$ if and only if $[A,L]=[A,[A,L]]$; see Section 2.
Following [PR04, Definition 5.1], a non-zero projection $p$ in a $C^{*}$-algebra is said to be weakly divisible of degree $n$ if there exists a unital *-homomorphism
$M_{n_{1}}({\mathbb{C}})\oplus M_{n_{2}}({\mathbb{C}})\oplus\ldots\oplus M_{n_{r}}({\mathbb{C}})\to pAp$
for some natural numbers $r\geq 1$ and $n_{1},\ldots,n_{r}\geq n$. The *-homomorphism may additionally be assumed to be injective.
As noted in [PR04, p.164], every matrix algebra $M_{n}({\mathbb{C}})$ for $n\geq 2$ admits a unital (not necessarily injective) *-homomorphism $M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})\to M_{n}({\mathbb{C}})$, and it follows that a projection $p$ weakly divisible of degree $2$ if and only if there exists a unital (not necessarily injective) *-homomorphism $M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})\to pAp$. We deduce that the unit in a $C^{*}$-algebra $A$ is weakly divisible of degree $2$ if and only if there exists an injective *-homomorphism
$M_{2}({\mathbb{C}})\to A,\quad M_{3}({\mathbb{C}})\to A,\quad\text{ or }\quad M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})\to A.$
In the first two cases, we get $A\cong M_{2}(B)$ and $A\cong M_{3}(C)$ for suitable unital $C^{*}$-algebras $B$ and $C$, and for such matrix algebras it was shown by Marcoux [Mar95, Theorem 2.6] that every Lie ideal is related to a two-sided ideal (which is also unique, as shown in [BKS08, Corollary 4.18]).
Our main contribution to Section 2 is therefore in the case that $A$ admits a unital (injective) *-homomorphism $M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})\to A$ but no unital *-homomorphism $M_{n}({\mathbb{C}})\to A$ for $n=2$ or $n=3$ (or actually any $n\geq 2$). The Cuntz algebra $\mathcal{O}_{\infty}$ is an important such example. (Using $K$-theory, we see that there exists no unital *-homomorphism $M_{n}({\mathbb{C}})\to\mathcal{O}_{\infty}$ for any $n\geq 2$. On the other hand, $\mathcal{O}_{\infty}$ admits a unital *-homomorphism from $M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})$ as explained in the proof of Section 5.)
Our approach is inspired by the methods developed by Marcoux [Mar95] to describe Lie ideals in matrix algebras, as well as the techniques of Brešar, Kissin, and Shulman [BKS08, Section 3] to study Lie ideals in algebras that decompose with respect to an idempotent. However, the result does not simply follow by combining these methods, for example by applying [BKS08, Theorem 3.4] for a suitable idempotent in the $C^{*}$-algebra. Specifically, given a unital *-homomorphism $M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})\to A$, we do not decompose $A$ with respect to the projection $p$ given by the image of the unit of $M_{2}({\mathbb{C}})$, in which case $pAp+(1-p)A(1-p)$ is a direct sum of matrix algebras (whence its Lie ideals are related to two-sided ideals), but it is unclear if (3.17) in [BKS08, Theorem 3.4] holds (or if $p$ is locally cyclic in the sense of [BKS08, Definition 3.5]). Instead, we use a ‘mixed’ decomposition with respect to a projection $g$ that is given by the image of the sum of a rank-one projection in $M_{2}({\mathbb{C}})$ with a rank-one projection in $M_{3}({\mathbb{C}})$, as explained in Section 2. In this case, one can show that $g$ is locally cyclic, which is the idea underlying the proof that $A_{12}LA_{12}\subseteq[A,[A,L]]$ in Lemma 2.1. On the other hand, the algebra $gAg+(1-g)A(1-g)$ is not a sum of matrix algebras and it is therefore not immediate that its Lie ideals are related to two-sided ideals.
[…]
Let $A$ be a unital $C^{*}$-algebra that admits a unital *-homomorphism
$\varphi\colon M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})\to A.$
Let $(e_{ij})_{i,j=1,2}$ denote the image of matrix units in $M_{2}({\mathbb{C}})$ under $\varphi$, and similarly let $(f_{ij})_{i,j=1,2,3}$ denote the image of matrix units in $M_{3}({\mathbb{C}})$ under $\varphi$. We do not assume that $\varphi$ is injective, and therefore the $e_{ij}$ or $f_{ij}$ may be zero.
Set
$g:=e_{11}+f_{11},\,\,\,\text{ and }\,\,\,h:=e_{22}+f_{22}+f_{33}.$
Then $1=g+h$, and we view elements in $A$ as operator matrices with respect to the decomposition induces by $g$ and $h$. More formally, we set
$A_{11}:=gAg,\quad A_{12}:=gAh,\quad A_{21}:=hAg,\,\,\,\text{ and }\,\,\,A_{22}:=hAh,$
and every element $a$ in $A$ can be written uniquely as a sum $a=a_{11}+a_{12}+a_{21}+a_{22}$ with $a_{ij}\in A_{ij}$ for $i,j=1,2$.
Lemma 2.1.
Retain the notation from Section 2, and let $L\subseteq A$ be a Lie ideal. Then the sets
$A_{11}LA_{22},\quad A_{22}LA_{11},\quad A_{12}LA_{12},\,\,\,\text{ and }\,\,\,A_{21}LA_{21}$
are all contained in $[A,[A,L]]$, and thus also in $[A,L]$ and in $L$.
Proof.
We verify that $A_{11}LA_{22}\subseteq[A,[A,L]]$. We first show that $gLh\subseteq L$. To see this, let $x\in L$. Then
$gx-xg=[g,x]\in[A,L],\,\,\,\text{ and }\,\,\,gx-2gxg+xg=[g,[g,x]]\in[A,[A,L]]\subseteq[A,L].$
Adding these expressions, we get
$2gxh=2gx-2gxg\in[A,L]\subseteq L,$
and using that $hg=0$ we get
$gxh=[g,gxh]=[\tfrac{1}{2}g,2gxh]\in[A,L]\subseteq L.$
We have shown that $gLh\subseteq L$.
Now, given $x\in L$ and $a,b\in A$, we have
$gagxhbh=[[gag,gxh],hbh]\in[[A,L],A]=[A,[A,L]].$
We have shown that $A_{11}LA_{22}\subseteq[A,[A,L]]$.
Analogously, one verifies that $A_{22}LA_{11}\subseteq[A,[A,L]]$.
We verify that $A_{12}LA_{12}\subseteq[A,[A,L]]$. Given $v\in A_{12}$ and $y\in L$, we have
(1) $vyv=[-\tfrac{1}{2}v,[v,y]]\in[A,[A,L]].$
Set
$v_{1}:=e_{12}+f_{12},\quad w_{1}:=e_{21}+f_{21},\quad v_{2}:=e_{12}+f_{13},\,\,\,\text{ and }\,\,\,w_{2}:=e_{21}+f_{31}.$
Then $v_{1},v_{2}\in A_{12}$ and $w_{1},w_{2}\in A_{21}$, and we have
$v_{1}w_{1}=v_{2}w_{2}=e_{11}+f_{11}=g,\quad w_{1}v_{1}=e_{22}+f_{22},\,\,\,\text{ and }\,\,\,w_{2}v_{2}=e_{22}+f_{33}.$
Now, given $x\in L$ and $a,b\in A_{12}=gAh$, set
$b^{\prime}=b(\tfrac{1}{2}e_{22}+f_{22}+f_{33}).$
Then $b^{\prime}$ belongs to $gAh$ and we have
$b=b^{\prime}(2e_{22}+f_{22}+f_{33}).$
We have already proved that $hAhLgAg\subseteq L$, which gives
$w_{1}axb^{\prime}w_{1}\in L,\,\,\,\text{ and }\,\,\,w_{2}axb^{\prime}w_{2}\in L.$
Applying (1) for $y=w_{1}axb^{\prime}w_{1}$ and $v=v_{1}$, we get
$axb^{\prime}(e_{22}+f_{22})=v_{1}(w_{1}axb^{\prime}w_{1})v_{1}\in[A,[A,L]].$
Analogously, we obtain that
$axb^{\prime}(e_{22}+f_{33})=v_{2}(w_{2}axb^{\prime}w_{2})v_{2}\in[A,[A,L]],$
and consequently
$axb=axb^{\prime}(2e_{22}+f_{22}+f_{33})\in[A,[A,L]].$
We have shown that $A_{12}LA_{12}\subseteq[A,[A,L]]$.
Analogously, one verifies that $A_{21}LA_{21}\subseteq[A,[A,L]]$. ∎
Lemma 2.2.
Retain the notation from Section 2, and let $L\subseteq A$ be a Lie ideal. Then
$AgLhA=AhLgA.$
Proof.
Using at the first step that $A=AgA=AhA$ since $g$ and $h$ are full, and applying at the third step that $hAgLhAg\subseteq L$ by Lemma 2.1, we get
$AgLhA=AhAgLhAgA=Ah(hAgLhAg)gA\subseteq AhLgA.$
Analogously, we have
$AhLgA=AgAhLgAhA=Ag(gAhLgAh)hA\subseteq AgLhA,$
as desired. ∎
Lemma 2.3.
Retain the notation from Section 2, and let $L\subseteq A$ be a Lie ideal. Then
$[A,L]\subseteq AgLhA.$
Proof.
To simplify notation, we set $I=AgLhA$. We have
$[A,gLh]\subseteq AgLh+gLhA\subseteq I.$
Further, using Lemma 2.2 at the third step, we have
$[A,hLg]\subseteq AhLg+hLgA\subseteq AhLgA=AgLhA=I.$
Thus, given $x\in L$, we have $[A,gxh+hxg]\subseteq I$. Note that $gxh,hxg\in L$ by Lemma 2.1. It follows that
$gxg+hxh=x-gxh-hxg\in L.$
We need to show that $[A,gxg+hxh]\subseteq I$. Using additivity of the Lie bracket and that $I$ is an additive subgroup, it suffices to verify that $[A_{ij},gxg+hxh]\subseteq I$ for $i,j=1,2$.
We show that $[A_{12},gxg+hxh]\subseteq I$. Given $a\in A$, we have
$[gah,gxg+hxh]=gahxh-gxgah.$
Using that $gxg+hxh\in L$, we have $[gah,gxg+hxh]\in L$, and thus
$[gah,gxg+hxh]\in L\cap gAh=gLh\subseteq AgLhA=I.$
We have shown that $[A_{12},gxg+hxh]\subseteq I$.
Similarly, and using also Lemma 2.2 at the last step, we get
$[hag,gxg+hxh]=hagxg-hxhag\in L\cap hAg=hLg\subseteq AhLgA=I$
for every $a\in A$. This shows that $[A_{21},gxg+hxh]\subseteq I$.
We show that $[A_{11},gxg+hxh]\subseteq I$. Given $a\in A$, set $y=gagxg-gxgag$. Then
$y=gagxg-gxgag=[gag,gxg+hxh]\in[A,L]\subseteq L.$
Then
$y(e_{12}+f_{12})=[y,e_{12}+f_{12}]\in[L,A]\subseteq L$
and thus
$y(e_{12}+f_{12})\in L\cap gAh=gLh\subseteq I.$
We deduce that
$y=y(e_{12}+f_{12})(e_{21}+f_{21})\in IA=I.$
We show that $[A_{22},gxg+hxh]\subseteq I$. Given $b\in A$, set $z=hbhxh-hxhbh$. Then
$z=hbhxh-hxhbh=[hbh,gxg+hxh]\in L\cap A_{22}.$
Then
$(e_{12}+f_{12})z=[e_{12}+f_{12},z]\in L\cap gAh\subseteq I,\,\,\,\text{ and }\,\,\,f_{13}z=[f_{13},z]\in I.$
We deduce that
$z=(e_{21}+f_{21})(e_{12}+f_{12})z+f_{31}f_{13}z\in AI+AI=I,$
as desired. ∎
[…]
Retain the notation from Section 2, and let $L\subseteq A$ be a Lie ideal. Then
$A[A,L]=A[A,L]A=AgLhA=AhLgA.$
Proof.
In general, if $M\subseteq A$ is a Lie ideal, then $AMA=AM$; see, for example, [GLT24, Lemma 3.2]. Since $[A,L]$ is a Lie ideal, we get $A[A,L]=A[A,L]A$.
By Lemma 2.3, we have $[A,L]\subseteq AgLhA$, and thus $A[A,L]A\subseteq AgLhA$. By Lemma 2.2, we have $AgLhA=AhLgA$. Finally, by Lemma 2.1, we have $gLh\subseteq[A,L]$, and thus $AgLhA\subseteq A[A,L]A$. ∎
Lemma 2.4.
Retain the notation from Section 2, and let $L\subseteq A$ be a Lie ideal, and set $I=A[A,L]A$. Then
$[A,I]\subseteq[A,[A,L]].$
Proof.
For each $i,j=1,2$, set $I_{ij}:=I\cap A_{ij}$. Since $I$ is a two-sided ideal, we have $I=I_{11}+I_{12}+I_{21}+I_{22}$. By Section 2, we have $I=AgLhA$, and thus
$I=AgLhA=\underbrace{gAgLhAg}_{I_{11}}+\underbrace{gAgLhAh}_{I_{12}}+\underbrace{hAgLhAg}_{I_{21}}+\underbrace{hAgLhAh}_{I_{22}}.$
By additivity of the Lie bracket and since $[A,L]$ is an additive subgroup, it suffices to show that $[A,I_{ij}]\subseteq[A,[A,L]]$ for each $i,j=1,2$.
By Lemma 2.1, we have
$I_{12}=gAgLhAh\subseteq[A,L],\,\,\,\text{ and }\,\,\,I_{21}=hAgLhAg\subseteq[A,L],$
and thus $[A,I_{12}]\subseteq[A,[A,L]]$ and $[A,I_{21}]\subseteq[A,[A,L]]$.
We show that $[A,I_{11}]\subseteq[A,[A,L]]$. It suffices to verify that $[A_{ij},I_{11}]\subseteq[A,[A,L]]$ for each $i,j=1,2$. Using that $hg=0$, and using Lemma 2.1 at the last step, we get
$[A_{12},I_{11}]=[gAh,gAgLhAg]=-(gAgLhAg)(gAh)=gAgLhAgAh\subseteq[A,[A,L]].$
Similarly, we have
$[A_{21},I_{11}]=[hAg,gAgLhAg]=(hAg)(gAgLhAg)=hAgAgLhAg\subseteq[A,[A,L]].$
We also have $[A_{22},I_{11}]=\{0\}\subseteq[A,[A,L]]$.
Further, using Lemma 2.1 at the last step, we have
$I_{11}=(gAgLh)Ag\subseteq[gAgLh,Ag]+Ag(gAgLh)\subseteq[A,L]+AgLh.$
Using again Lemma 2.1 at the last step, we get
$[A_{11},I_{11}]$ $\subseteq[A_{11},[A,L]]+[A_{11},AgLh]\subseteq[A,[A,L]]+[gAg,AgLh]$
$\subseteq[A,[A,L]]+gAgAgLh\subseteq[A,[A,L]].$
We have verified that $[A,I_{11}]\subseteq[A,[A,L]]$.
Analogously, one shows that $[A,I_{22}]\subseteq[A,[A,L]]$, which completes the proof. ∎
The next result verifies statement (1) in Section 1.
[…]
Let $A$ be a unital $C^{*}$-algebra that admits a unital *-homomorphism $M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})\to A$, and let $L\subseteq A$ be a (not necessarily closed) Lie ideal. Then, for the two-sided ideal $I:=A[A,L]A$, we have
(2) $[A,I]=[A,[A,L]]\subseteq[A,L]\subseteq I,\,\,\,\text{ and }\,\,\,[A,I]\subseteq L\subseteq T(I).$
Further, $I$ is the only two-sided ideal of $A$ to which $L$ is related.
Proof.
Using Lemma 2.3 at the first step, and Section 2 at the second step, we have
$[A,L]\subseteq AgLhA=I.$
This implies that $[A,[A,L]]\subseteq[A,I]$, and the converse inclusion $[A,I]\subseteq[A,[A,L]]$ is shown in Lemma 2.4. Using that $L$ is a Lie ideal, we have $[A,L]\subseteq L$, and therefore $[A,[A,L]]\subseteq[A,L]$. This shows the left chain of inclusions in (2).
The inclusion $[A,I]\subseteq L$ follows using that $[A,I]\subseteq[A,L]$ and $[A,L]\subseteq L$. Finally, since $[A,L]\subseteq I$, we have $L\subseteq T(I)$.
To show uniqueness of $I$, let $J$ be a two-sided ideal that is related to $L$, that is, such that
$[A,J]\subseteq L\subseteq T(J).$
Then $[A,L]\subseteq J$, and since $I$ is the two-sided ideal generated by $[A,L]$ we get $I\subseteq J$.
Applying Lemma 2.1 at the first step for $J$, considered as a Lie ideal, we get
$gJh\subseteq[A,J]\subseteq L,$
and thus
$gJh\subseteq L\cap gAh=gLh.$
Since $g$ and $h$ are full, we have $A=AgA$ and $A=AhA$, and consequently
$J=AJA=AgAJAhA=AgJhA\subseteq AgLhA=I.$
In conclusion, we get $I=J$, as desired. ∎
[…]
Let $A$ be a unital $C^{*}$-algebra that admits a unital *-homomorphism $M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})\to A$, and let $L\subseteq A$ be a closed Lie ideal. Then, the two-sided ideal $I:=A[A,L]A$ is closed.
Proof.
By Section 2, $L$ is related to $I$, that is, we have $[A,I]\subseteq L\subseteq T(I)$. Using that $L$ is closed, it follows that the closure $\overline{I}$ satisfies $[A,\overline{I}]\subseteq L\subseteq T(\overline{I})$. Since $I$ is the only two-sided ideal of $A$ to which $L$ is related (Section 2), we get $I=\overline{I}$. ∎
The next result verifies statement (2) in Section 1.
[…]
Let $A$ be a unital $C^{*}$-algebra that admits a unital *-homomorphism $M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})\to A$, and let $L\subseteq A$ be a (not necessarily closed) Lie ideal. Then the following statements are equivalent:
  1. (1)
    The Lie ideal $L$ is commutator equivalent to the two-sided ideal $A[A,L]A$.
  2. (2)
    The Lie ideal $L$ is embraced by some two-sided ideal.
  3. (3)
    We have $[A,L]=[A,[A,L]]$.
Proof.
It is clear that (1) implies (2). Assuming (2), let us show that (3) holds. Set $I:=A[A,L]A$, and let $J\subseteq A$ be some two-sided ideal that embraces $L$. Then $L$ is also related to $J$. It follows that $I=J$, since $I$ is the unique two-sided ideal to which $L$ is related, by Section 2. It follows that $L\subseteq T([A,I])$. Using this at the first step, and using Section 2 at the second step, we get
$[A,L]\subseteq[A,I]=[A,[A,L]],$
which verifies (3).
Assuming (3), let us show that (1) holds. Using that assumption at the firs step, and Section 2 at the second step, we get
$[A,L]=[A,[A,L]]=[A,I],$
which verifies (1). ∎
The next result is essentially contained in [PR04, Proposition 5.7]. We only show how to remove the separability assumption. A character on a $C^{*}$-algebra $A$ is one-dimensional, irreducible representation, that is, a surjective *-homomorphism $A\to{\mathbb{C}}$.
[…]
Let $A$ be a unital $C^{*}$-algebra of real rank zero. Then $A$ has no characters if and only if $A$ admits a unital *-homomorphism $M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})\to A$.
Proof.
The backwards implication is clear. To show the forward implication, assume that $A$ has no characters. By [KR15, Lemma 3.5], there exists a separable sub-$C^{*}$-algebra $A_{0}\subseteq A$ containing the unit of $A$. Using that real rank zero is separably inheritable (see [Bla06, Section II.8.5]), we obtain a separable sub-$C^{*}$-algebra $B\subseteq A$ of real rank zero with $A_{0}\subseteq B$. It follows that $B$ has no characters, which allows us to apply [PR04, Proposition 5.7] for $B$. We obtain a unital *-homomorphism from $M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})$ to $B$, and thus also to $A$. ∎
[…]
Let $A$ be a unital $C^{*}$-algebra of real rank zero that has no characters. Then Section 2 applies to $A$. In particular, every Lie ideal in $A$ is related to a (unique) two-sided ideal.
We expect that the following question has a positive answer.
[…]
Let $A$ be a unital $C^{*}$-algebra that has no characters. Is every Lie ideal in $A$ is related to a (unique) two-sided ideal?
3. Two-sided ideals in C*-algebras with weakly divisible unit
In this section, we prove that every two-sided ideal $I$ in a $C^{*}$-algebra $A$ that admits a unital *-homomorphism $M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})\to A$ satisfies $I=A[A,I]A$, and we describe $[A,I]$ as the subspace spanned by certain square-zero elements in $I$; see Section 3. If $I$ is a Dixmier ideal (in which case there is a well-defined ideal $I^{\frac{1}{2}}$ satisfying $(I^{\frac{1}{2}})^{2}=I$), then $I=[A,I]+[A,I^{\frac{1}{2}}]^{2}$; see Section 3. In this case every element in $I$ is a sum of square-zero elements in $I$ and of products of pairs of square-zero elements in $I^{\frac{1}{2}}$.
The next definition generalizes the notion of orthogonally factorizable square-zero elements in a ring ([GT23, Definition 5.1]) to a relative notion with respect to a two-sided ideal.
[…]
Let $I$ be a two-sided ideal in a $C^{*}$-algebra $A$. We let $N_{2}(I)$ denote the square-zero elements in $I$, that is,
$\operatorname{N}_{2}(I):=\big{\{}x\in I:x^{2}=0\big{\}}.$
We further set
$\operatorname{FN}_{2}(A,I):=\big{\{}x\in I:\text{ there exist }a,b\in A,y\in I\text{ with }x=ayb\mbox{ and }ba=0\big{\}}.$
Lemma 3.1.
Let $I\subseteq A$ be a two-sided ideal in a $C^{*}$-algebra, and let $x\in\operatorname{FN}_{2}(A,I)$. Then there exist $a,b\in A_{+}$ and $y\in\operatorname{FN}_{2}(A,I)\subseteq I$ such that
$x=ayb,\,\,\,\text{ and }\,\,\,ab=ba=0.$
Proof.
By definition, there are $c,d\in A$ and $z\in I$ such that
$x=czd,\,\,\,\text{ and }\,\,\,dc=0.$
Let $c=v|c|$ and $d=w|d|$ the polar decompositions in $A^{**}$ and set
$a:=|c^{*}|^{\frac{1}{2}},\quad b:=|d|^{\frac{1}{2}},\,\,\,\text{ and }\,\,\,y:=|c^{*}|^{\frac{1}{2}}vzw|d|^{\frac{1}{2}}.$
By properties of the polar decomposition (see, for example, [GKT23, Proposition 2.1]), we have $|c^{*}|^{\frac{1}{2}}v,w|d|^{\frac{1}{2}}\in A$, and thus $y\in A$. Further, we have $x=ayb$.
Since $dc=0$, we have $(d^{*}d)^{n}(cc^{*})^{m}=0$ for every $m,n\geq 1$. It follows that $p(d^{*}d)q(cc^{*})=0$ for every polynomials $p$ and $q$ with vanishing constant terms. Using functional calculus, we get $f(d^{*}d)g(cc^{*})=0$ for every continuous functions $f,g\colon\mathbb{R}\to\mathbb{R}$ with $f(0)=g(0)=0$. In particular, we get $|d|^{\frac{1}{2}}|c^{*}|^{\frac{1}{2}}=0$. Thus, we have $ba=0$, and then also $ab=0$.
To show that $y$ belongs to $\operatorname{FN}_{2}(A,I)$, consider the decomposition
$y=|c^{*}|^{\frac{1}{4}}\big{(}|c^{*}|^{\frac{1}{4}}vzw|d|^{\frac{1}{4}}\big{)}|d|^{\frac{1}{4}}.$
Then $|c^{*}|^{\frac{1}{4}}vzw|d|^{\frac{1}{4}}$ belongs to $I$, and we have also seen above that $|c^{*}|^{\frac{1}{4}}$ and $|d|^{\frac{1}{4}}$ have zero product, whence $y\in\operatorname{FN}_{2}(A,I)$. ∎
The next result is analogous to [GT23, Lemma 5.2].
Lemma 3.2.
Let $I\subseteq A$ be a two-sided ideal in a $C^{*}$-algebra. Then
$\operatorname{FN}_{2}(A,I)\subseteq[A,\operatorname{FN}_{2}(A,I)]\subseteq[A,[A,I]]\subseteq[A,I],\,\,\,\text{ and }\,\,\,\operatorname{FN}_{2}(A,I)\subseteq\operatorname{N}_{2}(I).$
Proof.
Let $x\in\operatorname{FN}_{2}(A,I)$. By Lemma 3.1, we can pick $a,b\in A_{+}$ and $y\in I$ such that $x=ayb$ and $ba=0$. As in the proof of Lemma 3.1, we see that $ba^{\frac{1}{2}}=0$. It follows that $a^{\frac{1}{2}}yb\in\operatorname{FN}_{2}(A,I)$ and then
$x=ayb=\big{[}a^{\frac{1}{2}},a^{\frac{1}{2}}yb\big{]}\in[A,\operatorname{FN}_{2}(A,I)].$
Since $\operatorname{FN}_{2}(A,I)$ is a subset of $I$, we have $[A,\operatorname{FN}_{2}(A,I)]\subseteq[A,I]$. Combining these results, and using at the last step that $[A,I]$ is a Lie ideal, we get
$\operatorname{FN}_{2}(A,I)\subseteq[A,\operatorname{FN}_{2}(A,I)]\subseteq[A,[A,\operatorname{FN}_{2}(A,I)]]\subseteq[A,[A,I]]\subseteq[A,I].$
Given $x=ayb\in\operatorname{FN}_{2}(A,I)$, with $a,b\in A$ and $y\in I$ such that $ba=0$, we have
$x^{2}=(ayb)(ayb)=0,$
and thus $x\in\operatorname{N}_{2}(I)$. ∎
[…]
We note that Section 3 and [GT23, Definition 5.1] agree when considering the ideal $I=A$ in a $C^{*}$-algebra $A$. Indeed, an element $x$ in $A$ belongs to $\operatorname{FN}_{2}(A)$ in the sense of [GT23, Definition 5.1] if there exist $s,t\in A$ such that $x=st$ and $ts=0$, while $x$ belongs to $\operatorname{FN}_{2}(A,A)$ as defined in Section 3 if there exist $a,y,b\in A$ such that $x=ayb$ and $ba=0$.
Given a factorization $x=ayb$ with $ba=0$, we can use $s=a$ and $t=yb$. Conversely, given a factorization $x=st$ with $ts=0$, it follows that $|t|^{\frac{1}{2}}s=0$. We consider the polar decomposition $t=v|t|$ in $A^{**}$. Then $v|t|^{\frac{1}{2}}$ belongs to $A$, and we can use $a=s$, $y=v|t|^{\frac{1}{2}}$ and $b=|t|^{\frac{1}{2}}$.
Given a subset $V$ in a complex vector space, we use $\operatorname{span}_{\mathbb{C}}V$ to denote the linear subspace generated by $V$. Following [GT23, Definition 2.6], we say that a ${\mathbb{C}}$-algebra $A$ is zero-product balanced if for all $a,b,c\in A$ we have
$ab\otimes c-a\otimes bc\in\operatorname{span}_{\mathbb{C}}\big{\{}u\otimes v\in A\otimes_{\mathbb{C}}A:uv=0\big{\}}\subseteq A\otimes_{\mathbb{C}}A.$
This is closely related to the concept of ‘zero-product determination’, for which we refer to [Bre21]. By [GT23, Corollary 2.12], a (not necessarily unital) $C^{*}$-algebra is zero-product balanced as a ${\mathbb{C}}$-algebra if and only if it is zero-product determined.
We see that the class of $C^{*}$-algebras considered in this section are zero-product balanced:
Lemma 3.3.
Let $A$ be a unital $C^{*}$-algebra that admits a unital *-homomorphism $M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})\to A$. Then $A$ is zero-product balanced as a ${\mathbb{C}}$-algebra.
Proof.
As in Section 2, let $(e_{ij})_{i,j=1,2}$ and $(f_{ij})_{i,j=1,2,3}$ be the images in $A$ of matrix units in $M_{2}({\mathbb{C}})$ and $M_{3}({\mathbb{C}})$. Then $e_{11}+f_{11}$ and $e_{22}+f_{22}$ are two full orthogonal projections in $A$. This allows us to apply [Rob16, Corollary 3.8], which shows that $A$ is generated by its projections as a ${\mathbb{C}}$-algebra, and thus is zero-product determined (hence, zero-product balanced) by [Bre21, Theorem 2.3]. ∎
A two-sided ideal $I\subseteq A$ in a $C^{*}$-algebra is said to be idempotent if $I=I^{2}$, that is, if every element in $I$ is of the form $x_{1}y_{1}+\ldots+x_{n}y_{n}$ for some $n\in{\mathbb{N}}$ and suitable $x_{j},y_{j}\in I$.
[…]
Let $A$ be a $C^{*}$-algebra that is zero-product balanced as a ${\mathbb{C}}$-algebra, and let $I\subseteq A$ be a two-sided ideal such that $I=AIA$ (for example, $A$ is unital or $I$ is idempotent). Then
$[A,I]=[A,[A,I]]=\operatorname{span}_{\mathbb{C}}\operatorname{FN}_{2}(A,I).$
Proof.
The inclusions $\supseteq$’ hold by Lemma 3.2. The proof of the converse inclusion is based on a combination of the methods used in the proofs of Proposition 2.14 and Theorem 5.3 in [GT23], which in turn are inspired by that of [Bre21, Theorem 9.1]. To verify that $[A,I]\subseteq\operatorname{span}_{\mathbb{C}}\operatorname{FN}_{2}(A,I)$, let $x\in A$ and $y\in I$. Set $F:=\operatorname{span}_{\mathbb{C}}\operatorname{FN}_{2}(A,I)$, and define
$\varphi_{y}\colon A\times A\to I/F,\quad\varphi_{y}(a,b):=bya+F.\qquad\qquad(a,b\in A)$
Note that $\varphi_{y}$ is a well-defined ${\mathbb{C}}$-bilinear map. Further, $\varphi_{y}$ preserves zero-products since if $a,b\in A$ satisfy $ab=0$, then $bya\in\operatorname{FN}_{2}(A,I)\subseteq F$. Using [GT23, Proposition 2.8] at the second step, we obtain that
$cyab+F=\varphi_{y}(ab,c)=\varphi_{y}(a,bc)=bcya+F$
and thus
$bcya-cyab\in F=\operatorname{span}_{\mathbb{C}}\operatorname{FN}_{2}(A,I)$
for all $a,b,c\in A$. The assumption that $I=AIA$ allows us to pick $n$ and $c_{j},a_{j}\in A$ and $y_{j}\in I$ for $j=1,\ldots,n$ such that $y=\sum_{j}c_{j}y_{j}a_{j}$. Then
$[x,y]=\sum_{j=1}^{n}[x,c_{j}y_{j}a_{j}]=\sum_{j=1}^{n}\big{(}xc_{j}y_{j}a_{j}-c_{j}y_{j}a_{j}x\big{)}\in\operatorname{span}_{\mathbb{C}}\operatorname{FN}_{2}(A,I),$
which implies that $[A,I]\subseteq\operatorname{span}_{\mathbb{C}}\operatorname{FN}_{2}(A,I)$. ∎
[…]
Does there exist a two-sided ideal $I$ in a $C^{*}$-algebra $A$ such that $[A,[A,I]]\neq[A,I]$?
The next result verifies statement (3) in Section 1.
[…]
Let $A$ be a unital $C^{*}$-algebra that admits a unital *-homomorphism $M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})\to A$, and let $I\subseteq A$ be a (not necessarily closed) two-sided ideal. Then
$I=A[A,I]A=A[A,I],\,\,\,\text{ and }\,\,\,[A,I]=[A,[A,I]]=\operatorname{span}_{\mathbb{C}}\operatorname{FN}_{2}(A,I).$
Proof.
Considering $I$ as a Lie ideal in $A$, it follows from Section 2 that $A[A,I]A$ is the unique two-sided ideal in $A$ to which $I$ is related. Since $I$ is clearly related to itself, we get $I=A[A,I]A$. The inclusion $A[A,I]\subseteq A[A,I]A$ holds since $A$ is unital. The converse inclusion holds since $[A,I]$ is a Lie ideal in $A$; see, for example, [GLT24, Lemma 3.2].
Now the equality $[A,I]=[A,[A,I]]$ also follows from Section 2. Further, since $A$ is zero-product balanced by Lemma 3.3, we have $[A,I]=\operatorname{span}_{\mathbb{C}}\operatorname{FN}_{2}(A,I)$ by Section 3. ∎
[…]
Let $A$ be a unital $C^{*}$-algebra that admits a unital *-homomorphism $M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})\to A$. Applying Section 3 for the ideal $A$, we get
$A=A[A,A]A,$
which means that $A$ is generated by its commutators as an ideal.
In fact, $A$ is even generated by its commutators as a ring. Indeed, by [GT25, Theorem 5.15], every element in $A$ is a sum of a commutator and the product of two commutators, and also the sum of three elements that are products of commutators.
The next result verifies statement (4) in Section 1.
[…]
Let $A$ be a unital $C^{*}$-algebra that admits a unital *-homomorphism $M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})\to A$. There is a natural bijection between the set $\mathcal{I}$ of two-sided ideals in $A$ and the set
:=$\mathcal{L}:=\big{\{}L\subseteq A\text{ Lie ideal}:L=[A,L]\big{\}}$
given by the maps $I\mapsto[A,I]$ and $L\mapsto A[A,L]A$.
Proof.
We denote the maps by $\alpha\colon\mathcal{I}\to\mathcal{L}$ and $\beta\colon\mathcal{L}\to\mathcal{I}$. Given $I\in\mathcal{I}$, using Section 3 at the second step, we have
$\alpha(I)=[A,I]=[A,[A,I]]=[A,\alpha(I)]$
which shows that $\alpha$ is well-defined.
Next, we show that $\alpha$ and $\beta$ are inverses of each other. First, given $I\in\mathcal{I}$, using Section 3 at the last step, we get
$\beta(\alpha(I))=\beta([A,I])=A[A,I]A=I.$
Second, given $L\in\mathcal{L}$, it follows from $L=[A,L]$ that $[A,L]=[A,[A,L]]$, and thus $L=[A,[A,L]]$. Using this at the last step, and using Section 2 at the third step, we get
$\alpha(\beta(L))=\alpha(A[A,L]A)=[A,A[A,L]A]=[A,[A,L]]=L,$
as desired, ∎
Robert showed in [Rob16, Lemma 2.1] that every nilpotent element in a $C^{*}$-algebra is a sum of commutators. With view towards Section 3, we ask:
[…]
Let $I$ be a (not necessarily closed) two-sided ideal in a $C^{*}$-algebra $A$. Does every nilpotent element in $I$ belong to $[A,I]$?
We give positive answers to Section 3 for arbitrary two-sided ideals in von Neumann algebras (Section 6), and for semiprime two-sided ideals in $C^{*}$-algebras (Section 4).
Following [GKT23, Definition 3.2], a Dixmier ideal in a $C^{*}$-algebra $A$ is a two-sided ideal $I\subseteq A$ such that $I$ is positively spanned (that is, $I=\operatorname{span}_{\mathbb{C}}(I\cap A_{+})$) and hereditary (that is, if $a,b\in A_{+}$ satisfy $a\leq b$ and $b\in I$, then $a\in I$) and strongly invariant (that is, if $x\in A$ satisfies $x^{*}x\in I$, then $xx^{*}\in I$). Two-sided ideals in von Neumann algebras are automatically Dixmier ideals, which was essentially shown by Dixmier (see [GKT23, Proposition 3.4]).
Given a Dixmier ideal $I$ and $s\in(0,\infty)$, there is a unique Dixmier ideal whose set of positive elements is $\{a^{s}:a\in I\cap A_{+}\}$, and we use $I^{s}$ to denote this ideal; see Proposition 3.7 and Definition 3.8 in [GKT23]. We note that for $n\in{\mathbb{N}}$ the ring-theoretic definition of $I^{n}$ as the two-sided ideal generated by $x_{1}\cdots x_{n}$ for $x_{1},\ldots,x_{n}\in I$ agrees with the ‘new’ definition of $I^{n}$ as the linear span of elements $a^{n}$ for $a\in I\cap A_{+}$. This leads to a well-behaved theory of roots and powers for Dixmier ideals ([GKT23, Theorem 3.9]), which we recall for the convenience of the reader.
[…]
Let $I\subseteq A$ be a Dixmier ideal in a $C^{*}$-algebra. Then
$(I^{s})^{t}=I^{st}=(I^{t})^{s},\,\,\,\text{ and }\,\,\,I^{s}I^{t}=I^{s+t}=I^{t}I^{s}$
for all $s,t\in(0,\infty)$.
Lemma 3.4.
Let $I\subseteq A$ be a Dixmier ideal in a $C^{*}$-algebra, let $x\in I$ with polar decomposition $x=v|x|$ in $A^{**}$, and let $s\in(0,\infty)$. Then
$v|x|^{s},|x|^{s},|x^{*}|^{s},|x|^{s}v^{*}\in I^{s}.$
Proof.
Since $I$ is a Dixmier ideal, we have $x^{*}\in I$, and thus $x^{*}x,xx^{*}\in I^{2}$. Then by Section 3, we get
$|x|=(x^{*}x)^{\frac{1}{2}}\in(I^{2})^{\frac{1}{2}}=I,\,\,\,\text{ and }\,\,\,|x^{*}|=(xx^{*})^{\frac{1}{2}}\in(I^{2})^{\frac{1}{2}}=I.$
It follows that $|x|^{s},|x^{*}|^{s}\in I^{s}$.
By [GKT23, Proposition 3.10], given a Dixmier ideal $J\subseteq A$, an element $y\in A$ belongs to $J^{\frac{1}{2}}$ if and only if $y^{*}y\in J$. Applying this for the Dixmier ideal $I^{2s}$, and using that
$\big{(}v|x|^{s}\big{)}^{*}\big{(}v|x|^{s}\big{)}=|x|^{2s}\in I^{2s},$
we deduce that $v|x|^{s}\in(I^{2s})^{\frac{1}{2}}=I^{s}$. Since $I^{s}$ is a Dixmier ideal, we further have $|x|^{s}v^{*}=(v|x|^{s})^{*}\in I^{s}$. ∎
The next result showcases a situation when every element in a Dixmier ideal $I$ is a sum of square-zero element in $I$ and products of pairs of square-zero elements in $I^{\frac{1}{2}}$. In Section 6, we strengthen this result for ideals in von Neumann algebras.
[…]
Let $A$ be a unital $C^{*}$-algebra that admits a unital *-homomorphism $M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})\to A$, and let $I\subseteq A$ be a Dixmier ideal. Then
$I=[A,I]+[A,I^{\frac{1}{2}}]^{2}.$
Proof.
We first show that
$A\cdot\operatorname{FN}_{2}(A,I)\subseteq[A,I]+[A,I^{\frac{1}{2}}]^{2}.$
Let $c\in A$ and $x\in\operatorname{FN}_{2}(A,I)$. By Lemma 3.1, we can pick $a,b\in A_{+}$ and $y\in I$ such that $x=ayb$ and $ab=ba=0$. Set $z:=yb\in I$, and let $z=v|z|$ be the polar decomposition in $A^{**}$. Then
$cx=\big{[}ca^{\frac{1}{2}},a^{\frac{1}{2}}z\big{]}+a^{\frac{1}{2}}zca^{\frac{1}{2}}=\big{[}ca^{\frac{1}{2}},a^{\frac{1}{2}}z\big{]}+\big{(}a^{\frac{1}{2}}v|z|^{\frac{1}{2}}\big{)}\big{(}|z|^{\frac{1}{2}}ca^{\frac{1}{2}}\big{)}.$
By Lemma 3.4, $v|z|^{\frac{1}{2}}$ and $|z|^{\frac{1}{2}}$ belong to $I^{\frac{1}{2}}$. Further, since $za=0$, we get $|z|^{\frac{1}{2}}a^{\frac{1}{2}}=0$, and hence
$a^{\frac{1}{2}}v|z|^{\frac{1}{2}}=\big{[}a^{\frac{1}{2}},v|z|^{\frac{1}{2}}\big{]}\in[A,I^{\frac{1}{2}}],\,\,\,\text{ and }\,\,\,|z|^{\frac{1}{2}}ca^{\frac{1}{2}}=\big{[}-ca^{\frac{1}{2}},|z|^{\frac{1}{2}}\big{]}\in[A,I^{\frac{1}{2}}].$
It follows that $cx$ belongs to $[A,I]+[A,I^{\frac{1}{2}}]^{2}$.
Using Section 3 at the first and second step, and using the above at the last step, we get
$I=A[A,I]=\operatorname{span}_{\mathbb{C}}\Big{(}A\cdot\operatorname{FN}_{2}(A,I)\Big{)}\subseteq[A,I]+[A,I^{\frac{1}{2}}]^{2}.$
On the other hand, we clearly have $[A,I]\subseteq I$. Using Section 3, we also have $[A,I^{\frac{1}{2}}]^{2}\subseteq(I^{\frac{1}{2}})^{2}=I$. ∎
4. Square-zero elements in semiprime ideals in C*-algebras
In this section, we study the subspace $N$ generated by the square-zero elements in a semiprime ideal in a $C^{*}$-algebra $A$. We show that $ANA$, the two-sided ideal generated by $N$, satisfies $ANA=[A,N]+N^{2}$; see Section 4. This generalize results from [GT24, Section 4].
We then characterize semiprimeness for two-sided ideals in a $C^{*}$-algebra whose unit is weakly divisible of degree $2$; see Section 4. In particular, such an ideal is semiprime if and only if it is generated by its commutators as an ideal (or as a ring).
We begin with some preparatory results about commutators involving Dixmier ideals, which might be of use in future work.
Lemma 4.1.
Let $I\subseteq A$ be a Dixmier ideal in a $C^{*}$-algebra. Then
$[A,I]\subseteq\big{[}I^{\frac{1}{2}},I^{\frac{1}{2}}\big{]}\subseteq I.$
Proof.
By Section 3, we have $I=I^{\frac{1}{2}}I^{\frac{1}{2}}$. Using that $[a,xy]=[ax,y]+[ya,x]$ for $a\in A$ and $x,y\in I$, it follows that
$[A,I]=\big{[}A,I^{\frac{1}{2}}I^{\frac{1}{2}}\big{]}\subseteq\big{[}AI^{\frac{1}{2}},I^{\frac{1}{2}}\big{]}+\big{[}I^{\frac{1}{2}}A,I^{\frac{1}{2}}\big{]}\subseteq\big{[}I^{\frac{1}{2}},I^{\frac{1}{2}}\big{]}.$
Further, we have $[I^{\frac{1}{2}},I^{\frac{1}{2}}]\subseteq I^{\frac{1}{2}}I^{\frac{1}{2}}=I$. ∎
Recall that $\operatorname{N}_{2}(I)$ denotes the set of square-zero elements in an ideal $I$.
Lemma 4.2.
Let $I\subseteq A$ be a Dixmier ideal in a $C^{*}$-algebra. Then
$\operatorname{N}_{2}(I)\subseteq\operatorname{span}_{\mathbb{C}}\big{[}\operatorname{N}_{2}(I^{\frac{1}{2}}),\operatorname{N}_{2}(I^{\frac{1}{2}})\big{]}.$
Proof.
Let $x\in\operatorname{N}_{2}(I)$, and let $x=v|x|$ be the polar decomposition in $A^{**}$. We have $x=|x^{*}|v=|x^{*}|^{\frac{1}{2}}v|x|^{\frac{1}{2}}$. Since $x^{2}=0$, the elements $|x|$ and $|x^{*}|$ are orthogonal, and hence so are $|x|^{\frac{1}{2}}$ and $|x^{*}|^{\frac{1}{2}}$. Therefore
$x=\big{[}\tfrac{1}{2}v|x|^{\frac{1}{2}},|x|^{\frac{1}{2}}-|x^{*}|^{\frac{1}{2}}\big{]}.$
By Lemma 3.4, the element $\tfrac{1}{2}v|x|^{\frac{1}{2}}$ belongs to $I^{\frac{1}{2}}$, and thus $\tfrac{1}{2}v|x|^{\frac{1}{2}}\in\operatorname{N}_{2}(I^{\frac{1}{2}})$. Set
$y:=|x|^{\frac{1}{2}}-|x^{*}|^{\frac{1}{2}}+v|x|^{\frac{1}{2}}-v^{*}|x^{*}|^{\frac{1}{2}}.$
It is elementary (but tedious) to check that $y^{2}=0$. Note that $y$ belongs to $I^{\frac{1}{2}}$ by Lemma 3.4, and thus $y\in\operatorname{N}_{2}(I^{\frac{1}{2}})$. Using that $v|x|^{\frac{1}{2}}$ and $v^{*}|x^{*}|^{\frac{1}{2}}$ also belong to $N_{2}(I^{\frac{1}{2}})$, it follows that
$|x|^{\frac{1}{2}}-|x^{*}|^{\frac{1}{2}}=y-v|x|^{\frac{1}{2}}+v^{*}|x^{*}|^{\frac{1}{2}}\in\operatorname{span}_{\mathbb{C}}N_{2}(I^{\frac{1}{2}}),$
and thus
$x=\big{[}\tfrac{1}{2}v|x|^{\frac{1}{2}},|x|^{\frac{1}{2}}-|x^{*}|^{\frac{1}{2}}\big{]}\in\operatorname{span}_{\mathbb{C}}\big{[}N_{2}(I^{\frac{1}{2}}),N_{2}(I^{\frac{1}{2}})\big{]},$
as desired. ∎
A two-sided ideal $I$ in a ring $R$ is said to be semiprime if for all two-sided ideals $J\subseteq R$ we have $J\subseteq I$ whenever $J^{2}\subseteq I$. We refer to [Lam01, Section 10] for details. In [GKT23, Section 5], together with Gardella and Kitamura, we showed that a two-sided ideal $I$ in a $C^{*}$-algebra is semiprime if and only if it is idempotent (that is, $I=I^{2}$), if and only if it is a Dixmier ideal and satisfies $I=I^{\frac{1}{2}}$.
The next result is a generalization of [GT24, Theorem 4.2] from the study of square-zero elements in $C^{*}$-algebras to square-zero elements in semiprime ideals in $C^{*}$-algebras. Recall that $\widetilde{A}$ denotes the minimal unitization of a $C^{*}$-algebra $A$.
[…]
Let $I\subseteq A$ be a semiprime ideal in a $C^{*}$-algebra, let $N$ denote the subspace generated by $\operatorname{N}_{2}(I)$, and let $J:=\widetilde{A}N\widetilde{A}$ denote the two-sided ideal of $A$ generated by $N$. Then $J$ is semiprime, and we have
$N=[N,N]\subseteq N^{2},\quad J=[J,J]^{2}=AN=[A,N]^{2}=[A,N]+N^{2},$
and
$[J,J]=[A,J]=[A,[A,N]]=[[J,J],[J,J]]=[[A,N],[A,N]].$
Proof.
We will use that $I$ is a Dixmier ideal satisfying $I=I^{\frac{1}{2}}$. Applying Lemma 4.2 at the second step, we have
$N=\operatorname{span}_{\mathbb{C}}\operatorname{N}_{2}(I)\subseteq\operatorname{span}_{\mathbb{C}}\big{[}\operatorname{N}_{2}(I^{\frac{1}{2}}),\operatorname{N}_{2}(I^{\frac{1}{2}})\big{]}=\operatorname{span}_{\mathbb{C}}\big{[}\operatorname{N}_{2}(I),\operatorname{N}_{2}(I)\big{]}=[N,N].$
The inclusion $[N,N]\subseteq N^{2}$ follows since every commutator in $[N,N]$ belongs to $N^{2}$. To verify that $[N,N]\subseteq N$, let $x,y\in\operatorname{N}_{2}(I)$. Then
$[x,y]=(1+x)y(1-x)-y+xyx.$
Each of the three summands is a square-zero element in $I$, which shows $[x,y]\in N$.
We have
$J:=\widetilde{A}N\widetilde{A}\subseteq\widetilde{A}N^{2}\widetilde{A}\subseteq\big{(}\widetilde{A}N\big{)}\big{(}N\widetilde{A}\big{)}\subseteq J^{2},$
which shows that $J$ is idempotent and therefore semiprime by [GKT23, Theorem A].
The inclusion $AN\subseteq\widetilde{A}N\widetilde{A}=J$ is clear. Conversely, given $a\in\widetilde{A}$, $x\in\operatorname{N}_{2}(I)$ and a unitary $u\in\widetilde{A}$, using that $u^{*}xu\in\operatorname{N}_{2}(I)$ and that $N\subseteq N^{2}$, we have
$axu=au(u^{*}xu)\in\widetilde{A}\operatorname{N}_{2}(I)\subseteq\widetilde{A}N\subseteq\widetilde{A}N^{2}\subseteq AN.$
Since every element in $\widetilde{A}$ is a linear combination of four unitaries, we get $J=AN$.
The inclusion $[A,N]+N^{2}\subseteq\widetilde{A}N\widetilde{A}=J$ is clear. For the converse inclusion, we use a similar argument as in the proof of Section 3 to show that $AN\subseteq[A,N]+N^{2}$. Given $a\in A$ and $x\in\operatorname{N}_{2}(I)$, let $x=v|x|$ be the polar decomposition in $A^{**}$. Since $I$ is semiprime, it is Dixmier and $I=I^{\frac{1}{2}}=I^{\frac{1}{4}}$. By Lemma 3.4, we have $v|x|^{\frac{1}{2}}\in I^{\frac{1}{2}}$ and thus $v|x|^{\frac{1}{2}}\in\operatorname{N}_{2}(I)$. Similarly, we have $v|x|^{\frac{1}{4}},|x|^{\frac{1}{4}}a|x^{*}|^{\frac{1}{2}}\in\operatorname{N}_{2}(I)$. Using that $x=|x^{*}|^{\frac{1}{2}}v|x|^{\frac{1}{2}}$, we get
$ax=\big{[}a|x^{*}|^{\frac{1}{2}},v|x|^{\frac{1}{2}}\big{]}+\big{(}v|x|^{\frac{1}{4}}\big{)}\big{(}|x|^{\frac{1}{4}}a|x^{*}|^{\frac{1}{2}}\big{)}\in[A,\operatorname{N}_{2}(I)]+\operatorname{N}_{2}(I)^{2}.$
Using that $J=AN$, we get $J=[A,N]+N^{2}$.
The inclusion $[A,N]^{2}\subseteq\widetilde{A}N\widetilde{A}=J$ is clear. Conversely, using that $N=[N,N]$, we have
$[A,N]=[A,[N,N]]\subseteq[N,[A,N]]\subseteq[[A,N],[A,N]]\subseteq[A,N]^{2}.$
We also have $N^{2}=[N,N]^{2}\subseteq[A,N]^{2}$ and thus
$J=[A,N]+N^{2}\subseteq[A,N]^{2}.$
The inclusion $[J,J]\subseteq[A,J]$ is clear. Using that $J=J^{2}$ and $[a,xy]=[ax,y]+[ya,x]$ for $a\in A$ and $x,y\in J$, we have
$[A,J]=[A,J^{2}]\subseteq[AJ,J]+[JA,J]\subseteq[J,J].$
This shows that $[J,J]=[A,J]$. We get
$[[A,N],[A,N]]\subseteq[A,[A,N]]\subseteq[A,J]=[J,J].$
We have
$\big{[}[A,N],N^{2}\big{]}\subseteq[A,N^{2}]\subseteq[AN,N]+[NA,N]\subseteq[A,N].$
and similarly $[N^{2},N^{2}]\subseteq[A,N^{2}]\subseteq[A,N]$. We also have
$[A,N]=[A,[N,N]]\subseteq[N,[A,N]]=[[N,N],[A,N]]\subseteq[[A,N],[A,N]].$
Using that $J=[A,N]+N^{2}$, it follows that
$[J,J]$ $\subseteq\big{[}[A,N],[A,N]\big{]}+\big{[}N^{2},[A,N]\big{]}+\big{[}[A,N],N^{2}\big{]}+\big{[}N^{2},N^{2}\big{]}$
$\subseteq\big{[}[A,N],[A,N]\big{]}.$
Using that $[A,N]\subseteq[A,J]=[J,J]$, we also have
$[J,J]=\big{[}[A,N],[A,N]\big{]}\subseteq\big{[}[J,J],[J,J]\big{]}.$
Since the converse inclusion is clear, we get
$[J,J]=[A,J]=[A,[A,N]]=[[J,J],[J,J]]=[[A,N],[A,N]].$
Finally, since $[A,N]\subseteq[J,J]$ and $J=[A,N]^{2}$, we have $J\subseteq[J,J]^{2}$ and thus $J=[J,J]^{2}$. ∎
Lemma 4.3.
Let $I\subseteq A$ be a Dixmier ideal in a $C^{*}$-algebra, and let $x\in I$ be nilpotent. Then $x\in[I^{\frac{1}{2}},I^{\frac{1}{2}}]$.
Proof.
As in the proof of [Rob16, Lemma 2.1], we proceed by induction on the degree of nilpotency. To start, let $x\in I$ be a square-zero element, and let $x=v|x|$ be the polar decomposition of $x$ in $A^{**}$. By Lemma 3.4, the elements $v|x|^{\frac{1}{2}}$ and $|x|^{\frac{1}{2}}$ belong to $I^{\frac{1}{2}}$, and thus
$x=\big{[}v|x|^{\frac{1}{2}},x^{\frac{1}{2}}\big{]}\in\big{[}I^{\frac{1}{2}},I^{\frac{1}{2}}\big{]}.$
Next, assume that for some $k\geq 2$ we have shown that every element $x\in I$ with $x^{k}=0$ belongs to $[I^{\frac{1}{2}},I^{\frac{1}{2}}]$. Let $x\in I$ satisfy $x^{k+1}=0$, and let $x=v|x|$ be the polar decomposition of $x$ in $A^{**}$. Set $y:=|x|^{\frac{1}{2}}v|x|^{\frac{1}{2}}$, which belongs to $A$. As in the proof of [Rob16, Lemma 2.1], one shows that $y^{k}=0$. By Lemma 3.4, we have $v|x|^{\frac{1}{2}},|x|^{\frac{1}{2}}\in I^{\frac{1}{2}}$. By assumption of the induction, we have $y\in[I^{\frac{1}{2}},I^{\frac{1}{2}}]$, and thus
$x=\big{[}v|x|^{\frac{1}{2}},|x|^{\frac{1}{2}}\big{]}+y\in\big{[}I^{\frac{1}{2}},I^{\frac{1}{2}}\big{]}$
as desired. ∎
[…]
Let $I\subseteq A$ be a semiprime ideal in a $C^{*}$-algebra, and let $x\in I$ be nilpotent. Then $x\in[A,I]$.
Proof.
Since $I$ is semiprime, it is a Dixmier ideal and satisfies $I=I^{\frac{1}{2}}$. Applying Lemma 4.3, we have
$x\in[I^{\frac{1}{2}},I^{\frac{1}{2}}]\subseteq[A,I^{\frac{1}{2}}]=[A,I],$
as desired. ∎
[…]
Let $A$ be a unital $C^{*}$-algebra that admits a unital *-homomorphism $M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})\to A$, and let $I\subseteq A$ be a two-sided ideal. The following statements are equivalent:
  1. (1)
    The ideal $I\subseteq A$ is semiprime.
  2. (2)
    We have $I=A[I,I]A$, that is, $I$ is generated by its commutators as an ideal in $A$.
  3. (3)
    We have $I=\widetilde{I}[I,I]\widetilde{I}$, that is, $I$ is generated by its commutators as an ideal.
  4. (4)
    We have $I=[I,I]^{2}$.
  5. (5)
    We have $[I,I]=[[I,I],[I,I]]$.
Proof.
To show that (2) implies (1), assume that $I$ is generated by $[I,I]$ as an ideal in $A$. We clearly have $[I,I]\subseteq I^{2}$, and therefore
$I=A[I,I]A\subseteq AI^{2}A=I^{2}.$
It follows that $I$ is idempotent, and thus semiprime by [GKT23, Theorem A].
It is clear that (4) implies (3), which in turn implies (2).
Assume that $I$ is semiprime, and let $N$ denote the subspace generated by the square-zero elements in $I$. Applying Section 3, we have
$I=A[A,I]A=\operatorname{span}_{\mathbb{C}}A\operatorname{FN}_{2}(A,I)A\subseteq ANA.$
Now it follows from Section 4 that
$I=ANA=[ANA,ANA]^{2}=[I,I]^{2},$
which verifies (4), and that
$[I,I]=[ANA,ANA]=[[ANA,ANA],[ANA,ANA]]=[[I,I],[I,I]],\,\,\,\text{ and }\,\,\,$
which verifies (5).
Finally, assuming (5) let us verify (1). Using that $[a,xy]=[ax,y]+[ya,x]$ for $a\in A$ and $x,y\in I$, we have
$[A,I^{2}]\subseteq[AI,I]+[IA,I]\subseteq[I,I].$
Applying Section 3 for the ideal $I^{2}$, we get
$I^{2}=A[A,I^{2}]A\subseteq A[I,I]A\subseteq A[[I,I],[I,I]]A\subseteq I^{4}.$
It follows that $I^{2}=I^{4}$, and thus semiprime by [GKT23, Theorem A]. ∎
5. Lie ideals in properly infinite C*-algebras
Let $A$ be a $C^{*}$-algebra with a unit that is weakly divisible of degree $2$, that is, that admits a unital *-homomorphism $M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})\to A$, and let $L\subseteq A$ be a (not necessarily closed) Lie ideal. We have seen in Section 2 that $L$ is related to the two-sided ideal $I:=A[A,L]A$. Further, by Section 2, $L$ is commutator equivalent to $I$ if and only if $[A,L]=[A,[A,L]]$.
This raises the question when Lie ideals $L$ in $A$ satisfy $[A,L]=[A,[A,L]]$. A sufficient condition is that $A=Z(A)+[A,A]$, where $Z(A)$ denotes the center of $A$; see Lemma 5.1. We observe in Section 5 that this condition is satisfies in a number of situations, in particular for all properly infinite $C^{*}$-algebras.
We then prove the main result of the paper: Every Lie ideal in a unital, properly infinite $C^{*}$-algebra is commutator equivalent to a (unique) two-sided ideal; see Section 5.
Lemma 5.1.
Let $A$ be a unital $C^{*}$-algebra such that $A=Z(A)+[A,A]$, and let $L\subseteq A$ be a Lie ideal. Then
$[A,L]=[A,[A,L]].$
Proof.
The inclusion $\supseteq$’ is clear. Conversely, using the assumption at the first step, and using the Jacobi identity at the last step, we get
$[A,L]=\big{[}Z(A)+[A,A],L\big{]}=\big{[}[A,A],L\big{]}\subseteq\big{[}A,[A,L]\big{]},$
as desired. ∎
Following Winter [Win12], a $C^{*}$-algebra $A$ is said to be pure if its Cuntz semigroup $\operatorname{Cu}(A)$ is almost unperforated and almost divisible. Pure $C^{*}$-algebras form a robust class [APTV24] that includes every $C^{*}$-algebra that tensorially absorbs the Jiang-Su algebra $\mathcal{Z}$, that is, $A\cong A\otimes\mathcal{Z}$. By [APT18, Theorem 7.3.11], a $C^{*}$-algebra $A$ is pure if and only if $\operatorname{Cu}(A)\cong\operatorname{Cu}(A)\otimes\operatorname{Cu}(\mathcal{Z})$.
By a quasitrace of a $C^{*}$-algebra $A$ we mean a lower-semicontinuous, $[0,\infty]$-valued $2$-quasitrace. By Haagerup’s theorem [Haa14], every bounded quasitrace on a unital, exact $C^{*}$-algebra is a trace, and this was later extended by Kirchberg [Kir97], who showed that arbitrary quasitraces on exact $C^{*}$-algebras are traces.
A unital $C^{*}$-algebra $A$ is weakly central if the center separates maximal ideals: Maximal ideals $M_{1},M_{2}\subseteq A$ satisfy $M_{1}=M_{2}$ whenever $M_{1}\cap Z(A)=M_{2}\cap Z(A)$. By Vesterstrøm’s theorem, $A$ is weakly central if and only if $A$ has the centre-quotient property: For every closed, two-sided ideal $I\subseteq A$, the quotient map $A\to A/I$ maps $Z(A)$ onto $Z(A/I)$.
A unital $C^{*}$-algebra $A$ has the Dixmier property if for every element $a\in A$ the closed convex hull of the unitary orbit $\{uau^{*}:u\in A\text{ unitary}\}$ has nonempty intersection with the center $Z(A)$. It was shown in [ART17, Theorem 1.1] that this property holds if and only if $A$ is weakly central, every simple quotient of $A$ has at most one tracial state, and every extreme tracial state of $A$ factors through a simple quotient. Dixmier showed that every von Neumann algebra has the property named after him.
[…]
Let $A$ be a unital $C^{*}$-algebra. Then $A=Z(A)+[A,A]$ in each of the following cases:
  1. (1)
    If $A$ is pure, every quasitrace on $A$ is a trace, and $A$ has the Dixmier property.
  2. (2)
    If $A$ has no tracial states (in particular, if $A$ is properly infinite).
  3. (3)
    If $A$ is a von Neumann algebra.
Proof.
(1) If $A$ has the Dixmier property, then every element in $A$ is a Dixmier element in the sense of [AGR23], and thus it follows from [AGR23, Lemma 4.10] that $A=Z(A)+\overline{[A,A]}$.
If $A$ is pure and every quasitrace on $A$ is a trace, then $[A,A]$ is closed by [NR16, Theorem 1.1]. Combining both results, we get $A=Z(A)+[A,A]$.
(2) By [Pop02, Theorem 1], we have $A=[A,A]$ if (and only if) $A$ has no tracial states.
(3) It was shown in the proof of [BKS08, Theorem 5.19] that every von Neumann algebra $A$ satisfies $A=Z(A)+[A,A]$. ∎
[…]
Let $A$ be a unital, properly infinite $C^{*}$-algebra, and let $L\subseteq A$ be a Lie ideal. Then for the two-sided ideal $I:=A[A,L]A$, we have
$[A,L]=[A,[A,L]]=[A,I].$
In particular, $L$ is commutator equivalent to $I$, and hence also embraced by $I$, and related to $I$.
Further, $I$ is the only two-sided ideal of $A$ to which $L$ is related.
Proof.
Since $A$ is unital and properly infinite, there exists a unital *-homomorphism from the Cuntz algebra $\mathcal{O}_{\infty}$ to $A$; see [Bla06, Proposition III.1.3.3]. Further, it is well-known that $\mathcal{O}_{\infty}$ admits a unital *-homomorphism from $M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})$, using for example Section 2 and that $\mathcal{O}_{\infty}$ is a unital $C^{*}$-algebra of real rank zero that admits no characters. It follows that $A$ admits a unital *-homomorphism $M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})\to A$.
We may therefore apply Section 2 and obtain that $L$ is related to $I$, and that $I$ is the only two-sided ideal to which $L$ is related. Further, since $A$ is unital and properly infinite, we have $A=[A,A]$ by Pop’s theorem [Pop02], and thus $[A,L]=[A,[A,L]]$ by Lemma 5.1. It therefore follows from Section 2 that $L$ is commutator equivalent to $I$ (and thus also embraced by $I$). ∎
[…]
Let $A$ be a unital, properly infinite $C^{*}$-algebra. Then an additive subgroup $L\subseteq A$ is a Lie ideal if and only if $[A,I]\subseteq L\subseteq T([A,I])$ for some (unique) two-sided ideal $I\subseteq A$. We get a decomposition of the family $\mathcal{L}$ of all Lie ideals in $A$ as:
$\mathcal{L}=\bigsqcup_{I\in\mathcal{I}}\big{\{}L\subseteq A\text{ additive subgroup}:[A,I]\subseteq L\subseteq T([A,I])\big{\}}$
where $\mathcal{I}$ denotes the lattice of two-sided ideals in $A$.
[…]
The proof of Section 5 can easily be adapted to show the following: In a $C^{*}$-algebra $A$ that admits a unital *-homomorphism $M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})\to A$ and such that $A=Z(A)+[A,A]$, every Lie ideal $L$ is commutator equivalent to the two-sided ideal $A[A,L]A$.
We note that a unital $C^{*}$-algebra admits no characters whenever it is pure or has no tracial stetes. Therefore, with view towards Section 2 and Section 5, we expect that every Lie ideal $L$ is commutator equivalent to $A[A,L]A$ whenever $A$ is a unital, pure, $C^{*}$-algebra with the Dixmier property and such that every quasitrace on $A$ is a trace, or whenever $A$ is unital and has no tracial states.
[…]
Given Lie ideals $L,M\subseteq A$ in a unital, properly infinite $C^{*}$-algebra, what is the relationship between the two-sided ideals generated by $[L,M]$, by $[A,[L,M]]$, and by $[A,L]$ and $[A,M]$?
6. Lie ideals and two-sided ideals in von Neumann algebras
In this section, we show that every (not necessarily closed) Lie ideal $L$ in a von Neumann algebra $M$ is commutator equivalent to the two-sided ideal $I:=M[M,L]M$, that is, $[M,L]=[M,I]$; see Section 6. This recovers [BKS08, Theorem 5.19]. Moreover, if $M$ has zero commutative summand, then we show that $I$ is the only two-sided ideal to which $L$ is related, which solves [BKS08, Problem 5.21].
Given a two-sided ideal $I$ in a von Neumann algebra with zero commutative summand, every element in $I$ is a sum of products of pairs of square-zero elements in $I^{\frac{1}{2}}$; see Section 6.
[…]
Let $M$ be a von Neumann algebra, and let $L\subseteq M$ be a Lie ideal. Then, for the two-sided ideal $I:=M[M,L]M$, we have
$[M,L]=[M,[M,L]]=[M,I].$
In particular, $L$ is commutator equivalent to $I$, and hence also embraced by $I$, and related to $I$.
If $M$ has zero commutative summand, then $I$ is the only two-sided ideal of $M$ to which $L$ is related.
Proof.
Let $M=M_{0}\oplus M_{1}$ be the (unique) decomposition of $M$ into a commutative summand $M_{0}$ and a summand $M_{1}$ that admits no characters. By Section 2, there is a unital *-homomorphism $M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})\to M_{1}$. Set
$L_{1}:=M_{1}\cap L,\,\,\,\text{ and }\,\,\,I_{1}:=M_{1}\cap I.$
Then $L_{1}$ is a Lie ideal in $M_{1}$, and we have
$[M,L]=[M_{1},L_{1}],\,\,\,\text{ and }\,\,\,I=I_{1}=M_{1}[M_{1},L_{1}]M_{1}.$
We may apply Section 2 for the Lie ideal $L_{1}$ in $M_{1}$ and obtain that $L_{1}$ is related to $I_{1}$, and that $I_{1}$ is the only two-sided ideal of $M_{1}$ to which $L_{1}$ is related. It follows that $L$ is related to $I$, and if $M_{0}$ is zero, then $I$ is also unique with this property. (If $M_{0}$ is nonzero, then $M_{0}\oplus I$ is another two-sided ideal of $M$ to which $L$ is related.)
As shown in the proof of [BKS08, Theorem 5.19], we have $M=Z(M)+[M,M]$. It follows that $[M,L]=[M,[M,L]]$ by Lemma 5.1, and then
$[M,L]=[M,[M,L]]=[M,I].$
by Section 2. ∎
Recall that $\operatorname{N}_{2}(I)$ denotes the square-zero elements, and we use $\operatorname{Nil}(I)$ to denote the set of nilpotent elements in an ideal $I$. Recall the definition of $\operatorname{FN}_{2}(M,I)$ from Section 3.
[…]
Let $I\subseteq M$ be a two-sided ideal in a von Neumann algebra. Then
$[M,I]=\operatorname{span}_{\mathbb{C}}\operatorname{N}_{2}(I)=\operatorname{span}_{\mathbb{C}}\operatorname{Nil}(I)=\operatorname{span}_{\mathbb{C}}\operatorname{FN}_{2}(M,I).$
Proof.
Let $M=M_{0}\oplus M_{1}$ denote the (unique) decomposition of $M$ into a commutative summand $M_{0}$ and a summand $M_{1}$ that admits no characters. Then
$[M,I]$ $=[M_{1},M_{1}\cap I],\quad\operatorname{N}_{2}(I)=\operatorname{N}_{2}(M_{1}\cap I),$
$\operatorname{Nil}(I)$ $=\operatorname{Nil}(M_{1}\cap I),\,\,\,\text{ and }\,\,\,\operatorname{FN}_{2}(M,I)=\operatorname{FN}_{2}(M_{1},M_{1}\cap I),$
which allows us to reduce the problem to studying the ideal $M_{1}\cap I$ in $M_{1}$.
Without loss of generality, we may therefore assume that $M$ has zero commutative summand. Then $M$ admits a unital *-homomorphism $M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})\to M$ by Section 2. Using Section 3 at the first step, we get
$[M,I]=\operatorname{span}_{\mathbb{C}}\operatorname{FN}_{2}(M,I)\subseteq\operatorname{span}_{\mathbb{C}}\operatorname{N}_{2}(I)\subseteq\operatorname{span}_{\mathbb{C}}\operatorname{Nil}(I).$
It remains to show that $\operatorname{Nil}(I)\subseteq[M,I]$. As in the proof of Lemma 4.3 (and [Rob16, Lemma 2.1]), we proceed by induction on the degree of nilpotency. To start, let $x\in I$ be a square-zero element. Let $x=v|x|$ be the polar decomposition in $M$. Then $|x|=v^{*}x\in I$. Since $x^{2}=0$, we have $|x|v=0$, and thus
$x=\big{[}v,|x|\big{]}\in[M,I].$
Next, assume that for some $k\geq 2$ we have shown that every element $x\in I$ with $x^{k}=0$ belongs to $[M,I]$. Let $x\in I$ satisfy $x^{k+1}=0$. Again, let $x=v|x|$ be the polar decomposition in $M$. We will show that the element $y:=|x|v$ satisfies $y^{k}=0$. Since $x^{k+1}=0$ and $|x|=v^{*}v|x|$, we have
$y^{k}|x|=|x|v\cdots|x|v|x|=v^{*}v|x|v\cdots|x|v|x|=v^{*}x^{k+1}=0.$
For every polynomial $p$ with vanishing constant term it follows that $y^{k}p(|x|)=0$. Using functional calculus, we get $y^{k}|x|^{\frac{1}{n}}$ for every $n\geq 1$. The sequence $(|x|^{\frac{1}{n}})_{n}$ converges in the weak*-topology of $M$ to $v^{*}v$, which is the support projection of $|x|$. We deduce that $y^{k}v^{*}v=0$. Using that $v=vv^{*}v$, we get
$y^{k}=y^{k-1}|x|v=y^{k-1}|x|vv^{*}v=y^{k}v^{*}v=0.$
By assumption of the induction, we have $y\in[M,I]$, and thus
$x=\big{[}v,|x|\big{]}+|x|v\in[M,I],$
as desired. ∎
[…]
Let $H$ be a separable, infinite-dimensional Hilbert space, and let $I$ be a two-sided ideal in the von Neumann algebra $\mathcal{B}(H)$. Given a nilpotent element $x\in I$, it follows from Section 6 that $x\in[\mathcal{B}(H),I]$, that is, $x$ is a finite sum of commutators of an element in $\mathcal{B}(H)$ and an element in $I$. Dykema and Krishnaswamy-Usha showed in [DKU18, Proposition 3.1] that one summand suffices, that is, we have $x=[y,z]$ for some $y\in\mathcal{B}(H)$ and $z\in I$. Does their result hold for nilpotent elements in two-sided ideals in arbitrary von Neumann algebras?
[…]
Let $M$ be a von Neumann algebra with zero commutative summand, and let $I\subseteq M$ be a two-sided ideal. Then
$I=\big{[}M,I^{\frac{1}{2}}\big{]}^{2}=\big{[}I^{\frac{1}{4}},I^{\frac{1}{4}}\big{]}^{2},$
and
$[M,I]\subseteq\big{[}[M,I^{\frac{1}{2}}],[M,I^{\frac{1}{2}}]\big{]}\subseteq\big{[}[I^{\frac{1}{4}},I^{\frac{1}{4}}],[I^{\frac{1}{4}},I^{\frac{1}{4}}]\big{]}\subseteq[I^{\frac{1}{2}},I^{\frac{1}{2}}].$
Proof.
We will use that $I$ is a Dixmier ideal, since two-sided ideals in von Neumann algebras are automatically Dixmier ideals ([GKT23, Proposition 3.4]). Using Section 6 at the first and last step, and using Lemma 4.2 at the second step, we have
$[M,I]=\operatorname{span}_{\mathbb{C}}\operatorname{N}_{2}(I)\subseteq\operatorname{span}_{\mathbb{C}}\big{[}\operatorname{N}_{2}(I^{\frac{1}{2}}),\operatorname{N}_{2}(I^{\frac{1}{2}})\big{]}=\big{[}[M,I^{\frac{1}{2}}],[M,I^{\frac{1}{2}}]\big{]}.$
Using Lemma 4.1, we then get
$[M,I]\subseteq\big{[}[M,I^{\frac{1}{2}}],[M,I^{\frac{1}{2}}]\big{]}\subseteq\big{[}[I^{\frac{1}{4}},I^{\frac{1}{4}}],[I^{\frac{1}{4}},I^{\frac{1}{4}}]\big{]}\subseteq\big{[}I^{\frac{1}{2}},I^{\frac{1}{2}}\big{]}$
By Section 2, there is a unital *-homomorphism $M_{2}({\mathbb{C}})\oplus M_{3}({\mathbb{C}})\to M$. We can therefore apply Section 3 at the first step, and together with the above we obtain that
$I=[M,I]+\big{[}M,I^{\frac{1}{2}}\big{]}^{2}\subseteq\big{[}M,I^{\frac{1}{2}}\big{]}^{2}.$
Applying Lemma 4.1 for the Dixmier ideal $I^{\frac{1}{2}}$, and using Section 3 at the last step, we get
$I\subseteq\big{[}M,I^{\frac{1}{2}}\big{]}^{2}\subseteq\big{[}I^{\frac{1}{4}},I^{\frac{1}{4}}\big{]}^{2}\subseteq\big{(}I^{\frac{1}{2}}\big{)}^{2}=I,$
as desired. ∎
[…]
It is likely that the inclusions involving $[M,I]$ in Section 6 are equalities. The key question is whether $[I^{\frac{1}{2}},I^{\frac{1}{2}}]$ is a subset of $[M,I]$. Using that $I$ is a Dixmier ideal and therefore admits a well-behaved theory of roots and powers (Section 3), we have
$\big{[}I^{\frac{1}{2}},I^{\frac{1}{2}}\big{]}=\big{[}I^{\frac{1}{4}}I^{\frac{1}{4}},I^{\frac{1}{2}}\big{]}\subseteq\big{[}I^{\frac{1}{4}},I^{\frac{1}{4}}I^{\frac{1}{2}}\big{]}+\big{[}I^{\frac{1}{2}},I^{\frac{1}{2}}I^{\frac{1}{4}}\big{]}=\big{[}I^{\frac{1}{4}},I^{\frac{3}{4}}\big{]},$
and by an analogous argument $[I^{\frac{1}{4}},I^{\frac{3}{4}}]\subseteq[I^{\frac{1}{8}},I^{\frac{7}{8}}]$. Inductively, it follows that $[I^{\frac{1}{2}},I^{\frac{1}{2}}]\subseteq[I^{\frac{1}{2^{n}}},I^{\frac{2^{n}-1}{2^{n}}}]$ for every $n\geq 1$, and thus
$\big{[}I^{\frac{1}{2}},I^{\frac{1}{2}}\big{]}\subseteq\bigcap_{n\geq 1}\big{[}I^{\frac{1}{2^{n}}},I^{\frac{2^{n}-1}{2^{n}}}\big{]}\subseteq\bigcap_{\varepsilon>0}\big{[}M,I^{1-\varepsilon}\big{]}.$
It is, however, not clear if $\bigcap_{\varepsilon>0}[M,I^{1-\varepsilon}]=[M,I]$.
For two-sided ideals $I,J\subseteq\mathcal{B}(H)$, it was shown in [DFWW04, Theorem 5.10] that $[\mathcal{B}(H),IJ]=[I,J]$. In particular, we have $[\mathcal{B}(H),I]=[I^{\frac{1}{2}},I^{\frac{1}{2}}]$, and we obtain the following consequence of Section 6:
[…]
Let $H$ be a separable, infinite-dimensional Hilbert space, and let $I$ be a two-sided ideal in $\mathcal{B}(H)$. Then
$[\mathcal{B}(H),I]=\big{[}[\mathcal{B}(H),I^{\frac{1}{2}}],[\mathcal{B}(H),I^{\frac{1}{2}}]\big{]}=\big{[}[I^{\frac{1}{4}},I^{\frac{1}{4}}],[I^{\frac{1}{4}},I^{\frac{1}{4}}]\big{]}=\big{[}I^{\frac{1}{2}},I^{\frac{1}{2}}\big{]}.$
[…]
Let $I,J\subseteq M$ be two-sided ideals in a von Neumann algebra. Do we have $[M,IJ]=[I,J]$?